### v3

# Injectivity radius lower bound of convex sum of tame Riemannian metrics and applications to symplectic topology

Abstract.
Motivated by the aspect of large-scale symplectic topology, we prove that for any pair $g_{0},\,g_{1}$ of smooth complete Riemannian metrics of bounded curvature and of injectivity radius bounded away from zero, the convex sum $g_{s}:=(1-s)g_{0}+sg_{1}$ also has bounded curvature depending only on the curvature bounds $\|R_{g_{i}}\|_{C^{0}}$ of $g_{0}$ or $g_{1}$, and that the injectivity radii of $g_{s}$ have uniform lower bound depending only on the derivative bounds $\|R_{g_{i}}\|_{C^{1}}=\|R_{g_{i}}\|_{C^{0}}+\|DR_{g_{i}}\|_{C^{0}}$. A main technical ingredient to establish the injectivity radius lower bound is an application of the quantitative inverse function theorem. Using these estimates, we prove that each quasi-isometry class of tame metrics is convex for all finite regularity class of $3\leq r<\infty$. Using this Riemannian geometry result, we prove that the set of smooth $C^{3}$-tame almost complex structures inside the same quasi-isometry class associated to the symplectic form $\omega$ is contractible in any $C^{k}$ topology for all finite $k\geq 0$.
Key words and phrases:
quasi-isometry, convex sum of metrics, geodesic flow, injectivity radii lower bounds, quantitative inverse function theorem, strong $C^{r}$ topology, $\mathfrak{T}$-tame almost complex structure, $C^{3}$-tameness of almost complex structure
2020 Mathematics Subject Classification:
53C20, 53D35
This work is supported by the IBS project # IBS-R003-D1.
1. Introduction
In a recent work of the present authors [CO24], we construct a Fukaya category on infinite-type surfaces and prove that the $A_{\infty}$ category is not quasi-equivalent to the colimit of Fukaya categories of finite-type surfaces. This category is an invariant under the deformation of tame $J$’s that is compatible to a fixed Riemannian metric of bounded curvature but not when the Riemannian metric goes out of the quasi-isometry class of the given metric. In this regard, the invariants arising from [CO24] (or any symplectic invariants relying on the structure of ideal boundary on noncompact symplectic manifolds in that matter) are not exactly symplectic invariants but are large-scale geometric symplectic invariants! This case concerns noncompact symplectic manifolds of infinite type.
In general higher dimensional situations, if a noncompact symplectic manifold $M$ has a suitably good compactification $\widehat{M}$ such as a complete Liouville manifold so that $\widehat{M}$ is given by the union
$M\sqcup\partial_{\infty}M$
together with the Liouville embedding $\partial_{\infty}M\times[0,\infty)\hookrightarrow M$, then the Liouville symplectic form $\omega=d\lambda$ carries a canonically associated quasi-isometry class $\mathfrak{T}$ induced by the image of $\partial_{\infty}M\times[0,\infty)$ in $M$. (See Remark 7.8 for the description of such a quasi-isometry class.) The growth rate invariants of symplectic homology of affine algebraic varieties is another such large-scale symplectic topological invariants. (See [Sei08], [McL12] for the study thereof.) This case concerns noncompact symplectic manifolds of finite type.
All objects of consideration of the present paper, such as Riemannian metrics and almost complex structures are assumed to be of $C^{\infty}$-class unless otherwise said.
1.1. Statements of main results: symplectic topology
Some more background and motivation of our study of large scale Riemannian geometry presented in the present article are now in order. The well-known Gromov’s lemma [Gro85] on noncompact symplectic manifolds holds in weak $C^{\infty}$ topology but fails to hold in strong $C^{\infty}$ topology if two metrics associated to the almost complex structures tame to a given symplectic form $\omega$ are allowed to vary beyond their quasi-isometry class, or not necessarily bilipschitz equivalent. (See Example 7.4 for such an example.) It is generally said that an almost complex structure $J$ on a symplectic manifold $(M,\omega)$ is called $\omega$-tame if the bilinear form $g_{J}:=\omega(\cdot,J\cdot)$ is a symmetric positive definite, i.e., defines a Riemannian metric and there exists a tame metric $g$ that satisfies (1.3) for some $A\geq 1$. Denote by ${\mathcal{J}}_{\omega}$ the set of such almost complex structures.
However to do the geometric analysis of pseudoholomorphic curves on noncompact symplectic manifolds, it is important to fix a tame behavior of the metric $g$ appearing here because the Riemannian metrics $g_{J_{0}}$ and $g_{J_{1}}$ associated to two $\omega$-compatible almost complex structures $J_{0}$ and $J_{1}$ are not necessarily quasi-isometric as mentioned above. When this happens, any symplectic invariants constructed using $J_{0}$ and $J_{1}$ via the pseudoholomorphic curves have no reason to be the same, when the construction involve the ideal boundary of the noncompact symplectic manifolds. The Fukaya category constructed for the infinite-type surface in [CO24] or the aforementioned growth rate of symplectic homology of affine algebraic varieties are examples of such invariants. A difference between the two situations is that the former concerns the case of non-cylindrical end while the latter does the case of cylindrical end.
The results on Riemannian geometry in the present paper and their implications to symplectic topology indicate that for the applications of the methodology of pseudoholomorphic curves to large scale symplectic topology of a noncompact symplectic manifold $(M,\omega)$, one needs to introduce the following notion of $(\omega,\mathfrak{T})$-tame almost complex structures, and assume at least $C^{3}$-tameness for the almost complex structures $J$ with respect to the symplectic form $\omega$. (See Definition 1.8 below for the definition of $C^{r}$-tameness.)
Definition 1.1 ($(\omega,\mathfrak{T})$-tame almost complex structures).
Let $\mathfrak{T}$ be a given quasi-isometry class of Riemannian metrics on $M$. We call $J$ a $C^{r}$ $(\omega,\mathfrak{T})$-tame almost complex structure if the following hold:
  1. (1)
    It is $\omega$-tame.
  2. (2)
    The Riemannian metric $g_{J}=\omega(\cdot,J\cdot)$ is in the quasi-isometry class $\mathfrak{T}$ and $C^{r}$-tame.
We denote by ${\mathcal{J}}_{\omega;\mathfrak{T}}$ the set of such almost complex structures.
Remark 1.2.
To the best knowledge of the present authors, such a $C^{3}$-requirement has not been recognized in the previous literature of symplectic topology. We believe that this $C^{3}$-tameness should be mentioned to ensure contractibiliy of $\omega$-tame almost complex structures with positive injectivity radius, and required in the definition of tame almost complex structures for the purpose of constructing (large scale) symplectic topological invariants of noncompact symplectic manifolds. This is because the natural candidate of the contraction to a reference tame metric $g_{\text{\rm ref}}$ given by
$(t,g)\mapsto tg_{\text{\rm ref}}+(1-t)g$ (1.1)
may not have a positive injectivity radius lower bound over $t\in[0,1]$ with $C^{r}$-tameness with $r<3$. See the next subsection, especially Remark 1.5, for further expounding of this Riemannian geometric difficulty.
Represent the given $\mathfrak{T}$ by a Riemannian metric $g$. A priori the answer to the question whether or not the subset ${\mathcal{J}}_{\omega;\mathfrak{T}}\subset{\mathcal{J}}_{\omega}$ is connected is not known because convexity of the set
$\{g_{J}\mid J\in{\mathcal{J}}_{\omega;\mathfrak{T}}\}=:\operatorname{Riem}_{\omega;\mathfrak{T}}(M)$ (1.2)
is not known as indicated in Remark 1.2 and [Nab96]. Recall that the ‘known proof’ of contractibility relies on the presumption that ‘the set of tame metrics is strictly convex’ so that the image of the path (1.1) is contained in (1.2).
In this regard we prove that the following result, the precise statement of which we refer to Theorem 8.2.
Theorem 1.3.
Let $(M,\omega)$ be a tame symplectic manifold. The set ${\mathcal{J}}_{\omega;\mathfrak{T}}$ consisting of smooth $C^{r}$-tame $r\geq 3$ is contractible with respect to the strong $C^{k}$ topology for $0\leq k<\infty$.
We refer readers to [Hir76] for the precise definition of strong $C^{r}$ topology which we briefly recall in the beginning of Section 5.
An immediate corollary of Theorem 1.3 is the following special case of two-dimensions. This is needed in the verification that the aforementioned Fukaya category constructed in [CO24] is indeed a large-scale symplectic topological invariant depending only on the quasi-isometry class of tame Riemannian metric. (See Theorem 1.9 stated below to see how the aspect of this Riemannian geometry enters.)
Corollary 1.4 (Theorem 10.3).
Let $\Sigma$ be a noncompact surface equipped with hyperbolic structure. Denote by $\mathfrak{T}$ a quasi-isometry class of hyperbolic structures of $\Sigma$ and by $\operatorname{Riem}_{\mathfrak{T}}(\Sigma)$ the set of Riemannian metrics quasi-isometric to the given hyperbolic metric. Then the set $\operatorname{Riem}_{\mathfrak{T}}(\Sigma)$ consisting of $C^{r}$-tame metrics of $3\leq r<\infty$ is contractible in strong $C^{k}$ topology for finite $0\leq k<\infty$.
The finding of the present article is a byproduct of the present authors’ work [CO24] on the Fukaya category of infinite-type surfaces in which their category strongly relies on the underlying hyperbolic structure of Riemann surfaces.
1.2. Statements of main results: Riemannian geometry
The standard proof of contractibility of the set of almost complex structures tame to a given symplectic form is based on the contractibility of relevant tame Riemannian metrics. The proof of the contractibility of tame Riemannian metrics is based on the convexity thereof.
Remark 1.5.
The main source of the requirement $r\geq 3$ for this $C^{r}$ tameness lies in the fact that the injectivity radius function $g\mapsto\iota_{g}$ is not a continuous function even with the strong $C^{\infty}$ topology on a noncompact manifold. Because of this, even when two complete Riemannian metrics $g_{0}$ and $g_{1}$ have positive injectivity radii $\iota_{g_{1}},\,\iota_{g_{2}}>0$, it is not a trivial matter to check in general, even with the $C^{\infty}$ topology, whether the positivity
$\inf_{t\in[0,1]}\{\iota_{g_{t}}\mid g_{t}:=(1-t)g_{0}+tg_{1}\}>0$
holds or not for the convex sum $t\mapsto(1-t)g_{0}+tg_{1}$. (See [Nab96] for the reason why.) The finding of the present paper shows that we need at least $C^{3}$-tameness for this positivity. At the moment it is not known whether this infimum is positive or not when $k\leq 2$.
To state the main results of the present paper on the global Riemannian geometry, we need to borrow the standard definitions of quasi-isometry and bilipschitz equivalence from large-scale geometry or coarse geometry.
Definition 1.6 (Quasi-isometry).
Two Riemannian metrics $g_{1},\,g_{2}$ on a smooth manifold are said to be quasi-isometric if there exists a constant $A\geq 1$ such that
$\frac{1}{A}\,g_{1}(u,u)\leq g_{2}(u,u)\leq A\,g_{1}(u,u)$ (1.3)
for all $u\in T_{x}M$ for all $x\in M$.
We also consider the following equivalence relation, which is a macroscopic version of the above quasi-isometry of Riemannian metrics.
Definition 1.7 (Bilipschitz).
Two Riemannian metrics $g_{1},\,g_{2}$ on a smooth manifold are said to be $A$-bilipschitz if their associated distance functions are $A$-bilipschitz, i.e., if there exists a constant $A\geq 1$ such that
$\frac{1}{A}d_{g_{1}}(x,y)\leq d_{g_{2}}(x,y)\leq Ad_{g_{1}}(x,y)$ (1.4)
for all $x,\,y\in M$. Here we denote by $d_{g_{i}}$ the distance function of $g_{i}$. We just say $g_{1},\,g_{2}$ are bilipschitz if they are $A$-bilipschitz for some $1\leq A<\infty$. We call the associated equivalence class a bilipschitz class.
It is easy to see that for the Riemannian distance metrics the above two notions of the above quasi-isometry and the bilipschitz equivalence are equivalent. (However one should recall that comparing two notions of ($C^{0}$) quasi-isometric equivalence and bilipshitz in general large-scale metric geometry or in coarse geometry is a highly nontrivial problem.)
We will denote by $\mathfrak{T}$ a quasi-isometry class or equivalently a bilipschitz class of a Riemannian metric. In the present paper, we will interchangeably use both terms as we feel more appropriate depending on the circumstances. We will also use the following terminology for the simplicity of exposition. We denote by $\operatorname{inj}_{g}:M\to{\mathbb{R}}_{+}$ the function of pointwise injectivity radius $x\mapsto\operatorname{inj}_{g}(x)$.
Definition 1.8.
Suppose that a complete Riemannian metric $g$ has positive injectivity radius lower bound $\iota_{g}:=\inf_{x\in M}\operatorname{inj}_{g}(x)>0$. We call a metric $g$ $C^{r}$-tame if there exists a constant $C_{\ell}>0$ for each $0\leq\ell\leq r-2$ such that
$\|D^{\ell}R_{g}\|_{C^{0}}=\sup_{x\in M}|D^{\ell}R_{g}(x)|<C_{\ell}.$
We just say that $g$ is tame if this holds for all $r\geq 2$.
From now on, we will always denote by $\mathfrak{T}$ a quasi-isometric class of tame Riemannian metrics, and by $\operatorname{Riem}_{\mathfrak{T}}(M)$ the set of tame metrics in class $\mathfrak{T}$, unless explicitly mentioned otherwise.
The first main result of Part I of the present paper is the following convexity result of $\mathfrak{T}$.
Theorem 1.9 (Theorem 3.5, 3.6 & 4.3).
[…]
  1. (1)
    There exists $\mathfrak{C}=\mathfrak{C}(C,\epsilon_{0},\epsilon_{1})>0$ such that
    $\|R_{s}\|_{C^{1}}\leq\mathfrak{C}.$ (1.5)
  2. (2)
    There exists some $\epsilon^{\prime}=\epsilon^{\prime}(C,\epsilon_{0},\epsilon_{1})>0$ such that
    $\iota_{g_{s}}\geq\epsilon^{\prime}>0$
    for all $s\in[0,1]$.
In particular, the set $\operatorname{Riem}_{\mathfrak{T}}(M)$ is strictly convex.
We would like to recall readers that estimating such a lower bound of the injectivity radius is a hard work as demonstrated by [Nab96], and involves certain volume control [Che70], [CGT82] in general. An upshot of the theorem is that the statement does not involve any volume control but that it crucially relies on the uniform bound for the $C^{2}$-norm of the exponential maps: This is one of the reasons why the derivative bound of the curvature, i.e., the $C^{3}$-tameness enters.
An outcome of these estimates, together with uniform curvature bound which will be also proved, enables us to prove that provided $3\leq r<\infty$,
  1. (1)
    the convex sum $s\mapsto(1-s)g_{0}+sg_{1}$ is contained in $\operatorname{Riem}_{\mathfrak{T}}(M)$,
  2. (2)
    it also defines a continuous path therein in strong $C^{r}$ topology.
Such a convexity of $\operatorname{Riem}_{\mathfrak{T}}(M)$ fails to hold for the strong $C^{\infty}$ topology in general, or under the lower regularity $1\leq r<3$. The latter is mainly because of the failure of Theorem 1.9 with that low regularity.
Theorem 1.10 (Contractibility in strong $C^{k}$ topology).
Let $\mathfrak{T}$ be any quasi-isometry class of $C^{r}$-tame Riemannian metrics on $M$ with $3\leq r<\infty$. Then $\operatorname{Riem}_{\mathfrak{T}}(M)$ is contractible in strong $C^{k}$ topology for all $0\leq k<\infty$.
In this regard, we do not know the answer to the following question, mainly because we do not have the uniform injectivity radii lower bound for the path under $C^{2}$ tameness.
Question 1.11.
Is the set $\operatorname{Riem}_{\mathfrak{T}}(M)$ strictly convex for a noncompact Riemannian manifold $M$, or is the path $s\mapsto(1-s)g_{0}+sg_{1}$ continuous in the subspace topology of $\operatorname{Riem}_{\mathfrak{T}}(M)$ in the strong $C^{2}$-topology of $\operatorname{Riem}(M)$?
We refer to Appendix B for some discussion on the contractibility with respect to some $C^{\infty}$ topology, which we call the direct limit $C^{\infty}$ topology.
1.3. Outline of the proofs
For the proof of uniform curvature bounds, we utilize the formula of the curvature operator $R_{g_{s}}$ of $g_{s}$ of convex sum $g_{s}:=(1-s)g_{0}+sg_{1}$ that is obtained in a recent article by Cavenaghi and Speranca [CS, Proposition 2.1].
The main part of the proof lies in that of the injectivity lower bound. The injectivity radii can jump under a continuous deformation even under the curvature bound because of the appearance or disappearance of a short geodesic which reflects a nonlocal behavior of metrics. By now it is well-known that uniformly controlling the injectivity radii even on compact manifolds is a difficult task in general. (We refer readers to [Nab96] to see why this is so.) Recall that a lower bound for the injectivity radius can be estimated by estimating the infimum of the radii of geodesic normal balls over the points of $M$. In particular we need to study the injectivity of the exponential maps to estimate the radii of such balls.
The main novelty of the present work is that the estimate of such radii under a continuation of metrics can be nicely done, although the estimate is not explicit, by exploiting the aforementioned bounds of the curvature and its derivative, and some idea of the proof of a quantitative inverse function theorem. For this purpose, we need the exponential maps to have uniform $C^{2}$-bounds. This is where our hypothesis on $C^{3}$-tameness of $g_{0}$ and $g_{1}$ enters. We would like to highlight that the proofs of the quantitative inverse function theorem, for example, from [AMR83, Corollary 2.5.6], [Chr85, Section 8] then enable us to rule out the appearance of short closed geodesics which is the main obstacle to have positive injectivity of radius bound, in the presence of a uniform derivative bound of the curvature.
To help readers get the overall scheme of our proof, we explain how each given condition put in the statement of Theorem 4.3 is used in the proof:
  1. (1)
    Injectivity radius bound of $g_{0}$ provides an atlas of $M$ with a uniform size of the coordinate geodesic normal balls of $g_{0}$.
  2. (2)
    Curvature bound $\|R_{g_{0}}\|_{C^{0}}$ and $\|R_{g_{1}}\|_{C^{0}}$ provides a uniform bound of $\|R_{g_{s}}\|_{C^{0}}$.
  3. (3)
    Quasi-isometry hypothesis on $g_{0}$ and $g_{1}$ and completeness thereof imply that $g_{s}$ are also complete and quasi-isometric thereto.
  4. (4)
    Completeness of $g_{s}$ and the bound for the derivatives $DR_{g_{0}}$ and $DR_{g_{1}}$ first implies the existence of a common domain and a codomain of the maps $F_{i}$ (see (4.15)), and then imply the uniform $C^{2}$ bound for the maps $F_{i}$.
  5. (5)
    Combining the above all, we can apply the quantitative inversion function theorem (Theorem 4.11) to conclude the uniform positive injectivity lower bound.
  6. (6)
    For the clarity of exposition, we consistently use the letter ‘r’ to talk about $C^{r}$-tameness, while we use the letter ‘k’ to talk about $C^{k}$-topology, respectively.
Once these basic estimates, especially the injectivity radius lower bound, are obtained, the proof of Theorem 5.1 will follow from the definition of strong $C^{k}$ topology of $\operatorname{Riem}_{\mathfrak{T}}(M)$, which is nothing but the subspace topology of the strong $C^{k}$ topology of $\operatorname{Riem}(M)$.
An interesting byproduct of this scheme of the proof is that any (finite-time) $C^{3}$-continuation of a tame metric on open manifold cannot develop a cusp, which seems to carry some interest of its own. Such a non-collapsing result under the uniform curvature bound in the study of the Ricci flow plays an important role for the application of Ricci flow to the 3-dimensional topology. (See [Per03], [MT07, Chapter 8].) In relation to the collapsing phenomenon under the Ricci flow, a finite-time extinction may be rephrased as a divergence in strong $C^{\infty}$ topology of the Riemannian metric under the Ricci flow in finite time.
Theorem 1.12.
Consider any continuous family $\{g_{s}\}_{s\in[0,1]}$ in strong $C^{3}$ topology of complete Riemannian metrics with
$\|R_{g_{0}}\|,\,\|DR_{g_{0}}\|_{C^{0}}<C,\quad\iota_{g_{0}}>0.$
Then there is a constant $C^{\prime}=C^{\prime}(C,\{g_{s}\}),\,\epsilon^{\prime}=\epsilon^{\prime}(C,\epsilon,\{g_{s}\})>0$ with $\epsilon:=\iota_{g_{0}}$ such that
$\inf_{s\in[0,1]}\iota_{g_{s}}>\epsilon^{\prime}$ (1.6)
Recalling that as the thin cylinder example [CGT82] or the cuspidal hyperbolic Riemann surfaces shows, the curvature bound itself does not provide the injectivity radius lower bound for a general single individual metric. The upshot of this theorem is that such a finite-time collapsing, or rather forming a cusp, cannot arise under a deformation of metrics that is continuous in strong $C^{3}$ topology.
Acknowledgement: We thank Gang Tian and Bruce Kleiner for useful email communications on the large-scale geometry, and the unknown referee for useful comments which helps us to improve the presentation of the paper.
Notations:
  1. (1)
    $\vec{a}$; a vector in ${\mathbb{R}}^{n}$,
  2. (2)
    $v$; an element in the tangent space $T_{p}M$,
  3. (3)
    $B^{n}(r)$; the standard open ball of radius $r$ centered at the origin of ${\mathbb{R}}^{n}$,
  4. (4)
    $B_{r}^{g}(p)$; the geodesic normal open ball of radius $r$ centered at $p\in M$,
  5. (5)
    $I_{p}^{g}:{\mathbb{R}}^{n}\to T_{p}M$; the canonical isometry with respect to a given orthonormal frame $\mathscr{B}$ of the inner product space $(T_{p}M,g_{p})$.
Part I Global Riemannian geometry
2. Convex sum of complete Riemannian metrics
Assume that both $g_{0}$ and $g_{1}$ are complete and of bounded geometry in the present section, and consider the convex sum thereof
$g_{s}=(1-s)g_{0}+sg_{1},\quad s\in[0,1].$
We will show that for any given reference tame metric $g_{\text{\rm rf}}$, the map
$(s,g)\mapsto(1-s)g+sg_{\text{\rm rf}}$
defines a contraction to the point $g_{\text{\rm rf}}$ that is continuous in strong $C^{k}$ topology of the set of $C^{r}$-tame metrics for $r\geq 3$. (It is not a priori contractible in the usual definition of strong $C^{\infty}$ topology. See Remark 5.2 for relevant comments.)
The main steps for the proof of Theorem 1.9 are the estimates of the curvatures and the injectivity radii bounds, especially the latter, for this convex combinations. Recently Cavenaghi and Speranca [CS] studied the curvature property of this convex sum in terms of the given metrics $g_{0},\,g_{1}$ for a different purpose. We use their explicit curvature formula to obtain some explicit bound for the curvature operator of $g_{s}$ in terms of those of $g_{0}$ and $g_{1}$.
We start with the following proposition on the completeness.
Proposition 2.1.
Suppose
$d_{0},\,d_{1}:{\mathcal{M}}\times{\mathcal{M}}\to{\mathbb{R}}_{\geq 0}$
are two complete metrics on the same topological space ${\mathcal{M}}$. If $d_{0},\,d_{1}$ are $A$-bilipschitz with $A\geq 1$ in the sense of Definition 1.7, $d_{s}$ is again $A$-bilipschitz with $d_{0},\,d_{1}$. In particular the space of $A$-bilipschitz complete metrics on ${\mathcal{M}}$ is convex.
Proof.
We first prove $A$-bilipschitz property of $d_{s}$. A direct calculation shows that $d_{s}$ satisfies
$\left((1-s)+\frac{s}{A}\right)d_{0}(x,y)\leq d_{s}(x,y)\leq((1-s)+sA)d_{0}(x,y)$
and
$\left(\frac{1-s}{A}+s\right)d_{1}(x,y)\leq d_{s}(x,y)\leq((1-s)A+s)d_{1}(x,y).$
Since $A\geq 1$, these inequalities show $d_{s}$ is $A$-bilipschitz with $d_{0},\,d_{1}$.
Completeness then immediately follows since the property is preserved under the bilipschitz equivalence. ∎
To prove the convexity of the bilipschitz class of tame metrics, we consider the convex sum of $g_{s}=(1-s)g_{0}+sg_{1}$ of $g_{0},\,g_{1}$. We have only to prove that $g_{s}$ is of bounded geometry, i.e.,
  1. (1)
    there exists a constant $C>0$ such that $\|R_{g_{s}}\|_{C^{0}}<C$ for the sectional curvature $R_{g_{s}}$.
  2. (2)
    there exists $\epsilon>0$ such that $\iota_{g_{s}}>\epsilon$ for all $s\in[0,1]$.
For the later purpose, we also consider the following $C^{0}$-distance between two metrics as a section of the bundle of symmetric positive definite quadratic forms.
Definition 2.2 (Quasi-isometric ratio).
Let $g_{0},\,g_{1}$ be two Riemannian metrics of $M$. We define
$A^{+}_{x}(g_{0},g_{1})$ ::=:= $\sup\{|v|_{g_{1}}\mid v\in T_{x}M,\,|v|_{g_{0}}=1\}$
$A^{-}_{x}(g_{0},g_{1})$ ::=:= $\inf\{|v|_{g_{1}}\mid v\in T_{x}M,\,|v|_{g_{0}}=1\}$
and
[…]
We call $A(g_{0},g_{1})$ the quasi-isometric ratio of $g_{0},\,g_{1}$.
By the definition of $A(g_{0},g_{1})$, $A(g_{0},g_{1})$ is finite if and only if $g_{0}$ and $g_{1}$ are quasi-isometric. It also follows that $A(g_{0},g_{1})=A(g_{1},g_{0})\geq 1$, and that $A(g_{0},g_{1})=1$ if and only if $g_{0}$ and $g_{1}$ are isometric.
Remark 2.3.
Consider the following standard notion in the comparison geometry: For a given pair of Riemannian metrics $g_{0},\,g_{1}$, we define the function $M_{g_{0},g_{1}}:M\to{\mathbb{R}}_{+}$ given by
$M_{g_{0},g_{1}}(x)=\sup_{0\neq v\in T_{x}M}\left|\log\left(\frac{|v|_{g_{1}}}{|v|_{g_{0}}}\right)\right|.$
(See Definition 7.5.) It follows that $M_{g_{0},g_{1}}$ is a continuous function on $M$. Then we have
$\log A(g_{0},g_{1})=\sup_{x\in M}M_{g_{0},g_{1}}(x).$
3. Bounds for the sectional curvature of the convex sum
We first consider the curvature estimate. We would like to recall readers that the curvature quadratically depends on the metric and its derivatives up to second order, and that considering the associated metric on the cotangent bundle amounts to taking the inverse of the metric coefficients and their derivatives. For our purpose, we will also need the bound for the derivative of the curvature and so will need the bound for the derivatives of the metric up to the third order.
This is the reason why getting the uniform bound for the curvature and its derivative of the convex sum $g_{s}$ in terms of the bounds for those of $g_{0},\,g_{1}$ needs to be verified, especially to obtain some explicit bound depending on the curvatures (and other tensorial expressions) and its derivatives of $g_{0},\,g_{1}$.
For this purpose, we follow the strategy used by Cavenaghi and Speranca [CS] in the following discussion. Consider the endomorphism $P=P_{g_{0};g_{1}}\in\operatorname{End}(TM)$ determined by
$g_{0}(PX,Y)=g_{1}(X,Y)$
which is positive definite symmetric with respect to $g_{0}$. We also consider another $D=D_{g_{0},g_{1}}\in\operatorname{End}(TM)$ given by the difference
$D_{g_{0},g_{1}}=\nabla^{1}-\nabla^{0}$
[…]
The following formula is proved in [CS].
Proposition 3.1 (Compare with Proposition 2.1 [CS]).
Let $R_{s}$ be the curvature operator associated to $g_{s}$. Then
$R_{s}(X,Y,Y,X)$ $=$ $(1-s)R_{0}(X,Y,Y,X)+sR_{1}(X,Y,Y,X)$ (3.1)
$+s(1-s)g_{1}\left((1-s)\text{\rm Id}+sP)^{-1}D(X,Y),D(Y,X)\right)$
$-s(1-s)g_{1}\left((1-s)\text{\rm Id}+sP)^{-1}D(Y,X),D(X,Y)\right).$
Here we mention that the endomorphism $(1-s)\text{\rm Id}+sP$ is invertible since $P$ is positive definite (with respect to $g_{0}$). We first state the following easy lemma.
Lemma 3.2.
Let $\lambda_{P}>0$ be the smallest eigenvalue of $P$. Then we have
$\left\|((1-s)\text{\rm Id}+sP)^{-1}\right\|_{C^{0}}\leq\frac{1}{1-s+s\lambda_{P}}\leq\frac{1}{\min\{1,\lambda_{P}\}}.$
From now on, we will take $g_{\text{\rm rf}}$ as the given reference metric which is also tame, and measure all relevant norms in terms of this metric $g_{\text{\rm rf}}$.
Now suppose the curvature bounds
$\|R_{g_{i}}\|_{C^{0}}\leq C_{i}<\infty$ (3.2)
[…]
Since $g_{0},\,g_{1}$ are bilipschitz and of bounded curvature, we also have the following estimate for the coordinate change matrices whose proof is essentially the same as that of [Che70, Lemma 3.4] and so omitted.
Lemma 3.3 (Compare with Lemma 3.4 [Che70]).
[…]
$\left|\frac{\partial y^{i}}{\partial x^{j}}(x)\right|<C(S)$ (3.3)
for all $x\in B_{0}\cap B_{1}$.
An immediate corollary of the curvature bound and this lemma is the following bound for the tensor $D$. This is a consequence of the general principle that curvature bounds imply bounds for the second derivatives of the metric tensor.
Proposition 3.4.
There exists some constant $C^{\prime}>0$ depending only on $g_{0},\,g_{1}$ (and $C_{0},\,C_{1}$) such that
$\|D\|_{C^{0}}\leq C^{\prime}$ (3.4)
Proof.
The bounds for the curvature and the injectivity radius lower bounds imply that we are given a pair of constants $r_{0},\,r_{1}$ given in Lemma 3.3 such that we can choose a normal coordinate system on a convex geodesic balls of the uniform size, say $\epsilon>0$, at every point $p\in M$ simultaneously for $g_{0},\,g_{1}$ so that
$B_{\epsilon}^{g_{\text{\rm rf}}}(p)\subset B_{r_{0}}^{g_{0}}(p)\cap B_{r_{1}}^{g_{1}}(p)$
and
$B_{\epsilon}^{g_{i}}(p)\subset B_{r_{0}}^{g_{0}}(p)\cap B_{r_{1}}^{g_{1}}(p).$
We fix such a constant $\epsilon>0$.
We first consider $g_{\text{\rm rf}}$ and denote the geodesic normal coordinates on $B_{\epsilon}^{g_{\text{\rm rf}}}(p)$ by $(x^{1},\cdots,x^{n})$ at $p\in M$ of $g_{0}$ By definition of normal coordinates, the associated normal coordinates $(x^{1},\cdots,x^{n})$ satisfies
$g_{ij}(p)=\delta_{ij},\quad\Gamma^{k}_{ij}(p)=0.$ (3.5)
Then we have the following Taylor expansion at $p$ of the metric tensor coefficients $g_{ij}$
$g_{ij}(x)$ $=$ $\delta_{ij}+\frac{1}{3}R_{ik\ell j}(p)x^{k}x^{\ell}+\frac{1}{6}R_{ik\ell j;s}(p)x^{k}x^{\ell}x^{s}$ (3.6)
$+\left(\frac{1}{20}R_{ik\ell j;st}(p)+\frac{2}{45}\sum_{m}R_{ik\ell m}(p)R_{jstm}(p)\right)x^{k}x^{\ell}x^{s}x^{t}$
$+O(r^{5})$
where $r$ is the distance from $p$. (See [MT07, Equation (1.5)] where the formula is attributed to Sakai [Sak96].)
We recall the formula for the Christoffel symbol
$\Gamma^{k}_{ij}=\frac{1}{2}g^{k\ell}\left(\frac{\partial g_{\ell j}}{\partial x^{i}}+\frac{\partial g_{i\ell}}{\partial x^{j}}-\frac{\partial g_{ij}}{\partial x^{\ell}}\right).$ (3.7)
By a direct calculation of derivatives of the right hand side of (3.6) at the given point $p$ of $g_{ij}$ using the expression (3.6) applied to $g_{0}$ and $g_{1}$ respectively;
$\frac{\partial g_{ij}}{\partial x^{m}}(x)=\frac{1}{3}\left(R_{imkj}(p)+R_{ikmj}(p)\right)x^{k}+O(r^{2})$
which implies
$\|\Gamma^{k}_{ij}(g_{0})\|_{C^{0};g_{0}}\leq C_{0}\|R_{g_{0}}\|_{C^{0};g_{0}}$ (3.8)
on $B_{r_{0}}^{g_{0}}(p)$. By the same token, we consider $g_{1}$ now and similarly have
$\|\Gamma^{k}_{ij}(g_{1})\|_{C^{0};g_{1}}\leq C_{1}\|R_{g_{1}}\|_{C^{0};g_{1}}$ (3.9)
on $B_{r_{1}}^{g_{1}}(p)$. Here the norms are measured by $g_{0}$ and $g_{1}$ respectively.
An immediate consequence of Lemma 3.3 is that the bounds (3.8), (3.9) can be converted to those in terms of the reference metric $g_{\text{\rm rf}}$,
[…]
by adjusting $C_{0},\,C_{1}$ slightly. We set $C^{\prime}=\max\{C_{0}^{\prime},C_{1}^{\prime}\}$ where we have
$C^{\prime}=C^{\prime}(C_{0},C_{1},r_{0},r_{1}).$
Then writing and substituting these estimates into the coordinate expression of $D=\nabla^{1}-\nabla^{2}$, we can find the bound $C_{k}^{\prime}$ depending only on
  • the curvature expression $R_{ik\ell j}^{g}(p)$ for $g=g_{0},\,g_{1}$ in coordinate functions $x^{i}$’s respectively,
  • the metric coefficients of $g_{0},\,g_{1}$ and their inverses.
These are valid on the geodesic ball
$B_{\epsilon}^{g_{\text{\rm rf}}}(p)\subset B_{r_{0}}^{g_{0}}(p)\cap B_{r_{1}}^{g_{1}}(p).$
Now we prove the curvature bound of $g_{s}$ in terms of that of $g_{0}$ and $g_{1}$ and the injectivity radius thereof.
Theorem 3.5.
Let $C_{0},\,C_{1}$ and $C^{\prime}=C^{\prime}(C_{0},C_{1},r_{0},r_{1})$ be as above. Then we have
$\|R_{s}\|_{C^{0}}\leq(1-s)C_{0}+sC_{1}+\frac{2s(1-s)}{\min\{1,\lambda_{P}\}}(C^{\prime})^{2}.$ (3.10)
Proof.
The proof immediately follows from expressing the formula (3.1) after writing the formula in coordinates on normal neighborhoods of the uniform size $\epsilon=\epsilon(r_{0},r_{1})>0$ and applying the estimates obtained from the above lemmata. ∎
By the same kind of reasoning after taking the derivatives of the identity (3.1), we also derive the following uniform $C^{1}$ bound.
Theorem 3.6.
Suppose $\max\{\|DR_{g_{0}}\|_{C^{0}},\|DR_{g_{1}}\|_{C^{0}}\}\leq C_{2}$ for some $C_{2}>0$ in addition. Then there exists some $C^{\prime\prime}=C^{\prime\prime}(C_{0},C_{1},r_{0},r_{1},C_{2})>0$ such that
$\max_{s\in[0,1]}\|DR_{s}\|_{C^{0}}\leq C^{\prime\prime}.$
4. Lower bounds for the injectivity radii of the convex sum
This is the central section of the present paper. The goal of this section is to prove a uniform lower estimate of injectivity radius for the convex sum. For this purpose, we will utilize some idea entering in the proof of a quantitative inverse function theorem. (See the proofs of [AMR83, Corollary 2.5.6] or of [Chr85, Section 8] which are given under the full uniform $C^{2}$ bounds on the maps.)
For each given tame metric $g$, we isometrically identify $(T_{p}M,g|_{p})$ with ${\mathbb{R}}^{n}$ with the standard inner product once and for all. We denote this identification map by
$I_{p}^{g}:{\mathbb{R}}^{n}\to T_{p}M$
by choosing an orthonormal frame $\mathscr{B}$ of $TM|_{B_{r_{0}}}^{g}(p)$ of $g$.
Remark 4.1.
Compare the map $I^{g}_{p}$ with the linear isometry, also denoted by $I:T_{\underline{m}}\underline{M}\to T_{m}M$, between two Riemannian manifolds $M,\,\underline{M}$ in [Che69]. The map is then composed with an exponential map to study some continuity property of the composition
$I\circ\exp_{\underline{m}}^{-1}:\underline{M}\to T_{m}M.$
One might regard the above map $I_{q}^{g}$ the same kind of map when $\underline{M}$ is the fixed universal space ${\mathbb{R}}^{n}$.
This map restricts to a diffeomorphism from
$U_{p}:=(I_{p}^{g})^{-1}\left((\exp_{p}^{g})^{-1}(B_{r_{0}}^{g}(p))\right)\subset{\mathbb{R}}^{n}$ (4.1)
to
$V_{p}:=(\exp_{p}^{g})^{-1}(B_{r_{0}}^{g}(p))\subset T_{p}M$ (4.2)
for each $p\in M$ provided $r_{0}<\iota_{g}$. We consider the map
$\psi_{q}^{g}=\exp_{q}^{g}\circ I_{q}^{g}$ (4.3)
at each point $q\in M$. The following is easy to check but is a key lemma that plays a fundamental role in our application of an argument of the quantitative inverse function theorem.
Lemma 4.2.
Suppose that the metric $g$ is complete. Then the maximal domain of the map $\psi_{q}^{g}$ is ${\mathbb{R}}^{n}$.
Proof.
This is an immediate consequence of the definition of $\psi_{q}^{g}=\exp_{q}^{g}\circ I_{q}^{g}$ from the completeness hypothesis of the metric so that the exponential map $\exp_{q}^{g}$ is globally defined at all $q\in M$. ∎
Let $g_{\text{\rm rf}}$ be any given reference tame metric and consider a constant
$0<r_{0}<\iota_{g_{\text{\rm rf}}}.$ (4.4)
We fix this $r_{0}$ and a coordinate atlas of $M$ given by
$\left\{\left((\psi_{q}^{g_{\text{\rm rf}}})^{-1},B^{g_{\text{\rm rf}}}_{r_{0}}(q)\right)\right\}_{q\in M};\quad\psi_{q}^{g_{\text{\rm rf}}}:=\exp_{q}^{g_{\text{\rm rf}}}\circ I_{q}^{g_{\text{\rm rf}}}$ (4.5)
which provides the Gaussian normal coordinates $(x^{1},\ldots,x^{n}):=(\psi_{q}^{g_{\text{\rm rf}}})^{-1}$ on the geodesic normal balls $B^{g_{\text{\rm rf}}}_{r_{0}}(q)$ associated to the metric $g_{\text{\rm rf}}$. Without loss of generality, we may assume that the geodesic normal ball
$B^{g_{\text{\rm rf}}}_{r_{0}}(q)=(\psi_{q}^{g_{\text{\rm rf}}})^{-1}(B^{n}(r_{0}))$
is strongly convex for all $q\in M$ by shrinking $r_{0}$ further if necessary. Such a shrinking of $r_{0}>0$ can be verified by the tame property of the reference metric $g_{\text{\rm rf}}$. (See [CE08, Appendix p.103] for its proof.)
For two different points $q,\,p$, we have the following diagram:
[…]
(We note that the map $\psi_{q}^{g}$ is globally defined on whole ${\mathbb{R}}^{n}$ as long as the metric $g$ is complete. Therefore the map restricts to an injective map on any open subset of the type $U_{q}$ defined above at every point $q\in M$ by the definition of geodesic normal ball $B_{r}^{g}(q)$.)
Recalling the definition of the injectivity radius function $\operatorname{inj}_{g}:M\to{\mathbb{R}}$ and the injectivity radius $\iota_{g}:=\inf_{x\in M}\operatorname{inj}_{g}(x)$, the following uniform lower estimate of the injectivity radii is the key result towards the proof of our convexity result.
Theorem 4.3.
[…]
$\iota_{g_{s}}\geq\epsilon^{\prime}>0$
for all $s\in[0,1]$.
As mentioned before, the main difficulty of the proof lies in the facts:
  1. (1)
    The set of $C^{r}$ tame metrics may not be convex with low regularity $1\leq r<3$.
  2. (2)
    The function $s\mapsto\iota_{g_{s}}$ is not a continuous function in strong $C^{k}$ topology.
Our strategy of proving the theorem is as follows: Let $B\subset[0,1]$ be the set of $s\in[0,1]$ such that $\iota_{g_{s}}>0$. By the given hypothesis, $B$ is nonempty. Then we will show that $B$ is open and closed which will then show that $B=[0,1]$ by the connectedness of the interval [0,1][0,1][ 0 , 1 ]. After that, we will find a lower bound $\epsilon^{\prime}>0$ of injectivity radius and finish the proof.
4.1. Implication of curvature estimates: openness
Let $s\in[0,1]$ for which $\iota_{g_{s}}>0$. We will find some $\delta>0$ and $\epsilon>0$ for which $\iota_{g_{s^{\prime}}}>\epsilon$ for all $s^{\prime}\in(s-\delta,s+\delta)\cap[0,1]$. We consider the initial value problem for the geodesic equation of the metric $g_{s^{\prime}}$
$\nabla_{\dot{\gamma}}^{g_{s^{\prime}}}\dot{\gamma}=0,\quad\gamma(0)=p$ (4.7)
in terms of the Gaussian normal coordinates $(x^{1},\cdots,x^{n})$ at each $p$ and $s\in[0,1]$. With respect to the associated canonical coordinates
$(x^{1},\cdots,x^{n},v^{1},\cdots,v^{n})$
of $(x^{1},\cdots,x^{n})$, the equation becomes the equation of geodesic flow of $g_{s}$ given by
$\begin{cases}\dot{x}^{j}=v^{j},\\ \dot{v}^{j}=-\Gamma^{j}_{k\ell}(x^{1},\cdots,x^{n})v^{k}v^{\ell}.\end{cases}$ (4.8)
We recall that the geodesic flow is the Hamiltonian flow of the kinetic energy of $g_{s}$ on $TM$ which is globally defined on $TM$, provided the metric $g$ is complete. The equation (4.8) is the coordinate expression thereof in the canonical coordinates. (See [Kli78] for detailed explanations of this point of view.)
Remark 4.4.
It follows from the expression (3.7) of the Christoffel symbols that the local existence, uniqueness and continuity of solutions of (4.8) hold as long as the metric $g$ is in the class of $C^{1,1}$ and uniform if there is a uniform bound on $C^{1,1}$-norm as in the case when there are uniform bounds on the curvature and the quasi-isometric ratio $A(g,g_{\text{\rm rf}})$ with respect to a given back-ground metric $g_{\text{\rm rf}}$.
The following lemma is standard, which states that the geodesic flow on $TM$ is nothing but the Hamiltonian flow of $K_{g}$ for any Riemannian metric $g$.
Lemma 4.5.
Let $\phi_{K_{g}}^{t}$ be the flow of the vector field $X_{K_{g}}$ and $\pi:TM\to M$ the canonical projection. Then
$\exp_{p}^{g}(v)=\pi\circ\phi_{K_{g}}^{1}(p,v)$
and the metric $g$ is complete if and only if the vector field $X_{K_{g}}$ is complete.
We apply this lemma to the present one-parameter family $\{g_{s}\}$ of metrics $g_{s}$. We denote by $K_{s}:TM\to{\mathbb{R}}$ the kinetic energy Hamiltonian of $g_{s}$ with $s\in[0,1]$, $X_{K_{s}}$ the associated Hamiltonian vector field and $\phi_{K_{s}}^{t}$ the associated Hamiltonian flow which is nothing but the geodesic flow of the metric $g_{s}$. Since $g_{s}$ is complete, the Hamiltonian flow thereof is well-defined for all time $t\in{\mathbb{R}}$, in particular a complete trajectory of the geodesic flow of each $g_{s}$ with the initial condition $(x(0),v(0))=(p,v_{0})$ exists for every $v_{0}\in T_{p}M$.
Furthermore the curvature bound implies the following bounds.
Lemma 4.6.
Under the same hypotheses as above in the standing hypotheses of Section 3, there is a uniform constant $C^{\prime\prime\prime}>0$ for which on $B_{r_{0}}^{g_{\text{\rm rf}}}(q)$
$\|\Gamma^{j}_{k\ell}\|_{C^{0}},\quad\left\|\Gamma^{j}_{k\ell;m}\right\|_{C^{0}}\leq C^{\prime\prime\prime}$ (4.9)
holds for all $j,\,k,\,\ell,\,m$.
Proof.
Since the first bound is already proved in the course of the proof of Proposition 3.4, we have only to prove the second bound. For this, we differentiate (3.6) twice and obtain
$g_{ij;km}(p)=\frac{1}{3}R_{ikmj}(p)$
in any geodesic normal coordinates at each $p$. This provides the bound
$\|g_{ij;km}\|_{C^{0}}<C\quad\text{\rm on }\,B_{r_{0}}^{g_{i}}(p)$
[…]
Therefore since $\Gamma_{ij;m}$ is a sum of polynomials of $g_{ij;km}$, $g_{ij}$ and $g^{ij}$ of degree at most 3, we have finished the proof again by applying Lemma 3.3. ∎
We now consider the parameterized exponential map
$(s,t,v)\mapsto\exp_{p}^{g_{s}}(tv)=\pi\circ\phi_{K_{s}}^{t}(p,v).$ (4.10)
By definition, we have the (space) tangent map of $\exp_{p}^{g_{s}}:T_{p}M\to M$
$d\exp_{p}^{g_{s}}=d\pi d\phi_{K_{s}}^{t}$
for each $s$. We recall $d\exp_{p}^{g_{s}}(0):T_{p}M\to T_{p}M$ is the identity map for all $s\in[0,1]$.
Proposition 4.7.
Suppose $\iota_{g_{s}}>0$ for an $s\in[0,1]$. There exists a sufficiently small $\delta,\,\epsilon>0$ such that
$\iota_{g_{s^{\prime}}}>\epsilon$
for all $s^{\prime}\in(s-\delta,s+\delta)\cap[0,1]$.
Proof.
By the standing hypothesis, at each point $p\in M$, there exists an atlas of the form
$\left\{\left((\psi_{p}^{g_{s}})^{-1},B_{\iota_{g_{s}}}^{g_{s}}(p)\right)\right\}_{p\in M}.$
We denote by
$(x_{1},\cdots,x_{n})=(\psi_{p}^{g_{s}})^{-1}$
the associated geodesic normal coordinates of $g_{s}$ on $B_{\iota_{g_{s}}}^{g_{s}}(p)$. For any $s^{\prime}\in[0,1]$, we may apply Lemma 3.3 and get $\kappa_{s^{\prime}}>0$ so that
$B_{\kappa_{s^{\prime}}}^{g_{s^{\prime}}}(p)\subset B_{\iota_{g_{s}}}^{g_{s}}(p).$
Now let $r_{s}:=\inf_{s^{\prime}\in[0,1]}\kappa_{s^{\prime}}$. Then $r_{s}>0$ and $\exp_{p}^{g_{s}}\circ I_{p}^{g_{s}}$ is a diffeomorphism of $B^{n}(\iota_{g_{s}})$ onto its image $B_{\iota_{g_{s}}}^{g_{s}}(p)$ in $M$. Also, the inclusion
$(\exp_{p}^{g_{s^{\prime}}}\circ I_{p}^{g_{s^{\prime}}})(B^{n}(r_{s}))\subset(\exp_{p}^{g_{s}}\circ I_{p}^{g_{s}})(B^{n}(\iota_{g_{s}}))$
implies that the composition
$(\exp_{p}^{g_{s}}\circ I_{p}^{g_{s}})^{-1}\circ(\exp_{p}^{g_{s^{\prime}}}\circ I_{p}^{g_{s^{\prime}}}):B^{n}(r_{s})\subset{\mathbb{R}}^{n}\to B^{n}(\iota_{g_{s}})\subset{\mathbb{R}}^{n}$
is defined for all $s^{\prime}\in[0,1]$ by the completeness of $g_{s^{\prime}}$. The latter also implies the exponential map $\exp_{p}^{g_{s^{\prime}}}$ is defined on whole $T_{p}M$.
Lemma 4.8.
Let $s$ be given and let $r_{s}>0$ be as above. There exists a sufficiently small $\epsilon_{s}>0$ and in turn a $\delta>0$ such that for all $s^{\prime}\in(s-\delta,s+\delta)\cap[0,1]$
$\|(\psi_{p}^{g_{s}})^{-1}\circ\psi_{p}^{g_{s^{\prime}}}-\operatorname{Id}\|_{C^{1}}<\epsilon_{s}$
on $B^{n}(r_{s})$ where $\epsilon_{s}$ can be made as small as we want by taking $\delta>0$ small.
Proof.
We mention that $g_{s^{\prime}}\to g_{s}$ in local $C^{\infty}$ topology as $s^{\prime}\to s$. In particular
$(\psi_{p}^{g_{s}})^{-1}\circ\psi_{p}^{g_{s^{\prime}}}\to\operatorname{Id}$
on $\overline{B}^{n}(r_{s})$ in $C^{\infty}$ topology and so in $C^{1}$-topology. The lemma follows from this. ∎
In particular the map
$(\psi_{p}^{g_{s}})^{-1}\circ\psi_{p}^{g_{s^{\prime}}}=:\Phi_{ss^{\prime}}$
is a homeomorphism (and so a diffeomorphism) from $B^{n}(r_{s})$ onto its image $\Phi_{ss^{\prime}}(B^{n}(r_{s}))$. By letting $\delta>0$ smaller if necessary, we may assume that
$\overline{B^{n}(r_{s^{\prime}})}\subset\Phi_{ss^{\prime}}(B^{n}(r_{s}))$ (4.11)
where we can choose $r_{s^{\prime}}=\frac{r_{s}}{2}$ for every $s^{\prime}\in(s-\delta,s+\delta)\cap[0,1]$ which does not depend on $s^{\prime}$.
Then the map
$(\psi_{p}^{g_{s^{\prime}}})^{-1}$ $=$ $(\exp_{p}^{g_{s^{\prime}}}\circ I_{p}^{g_{s^{\prime}}})^{-1}=(I_{p}^{g_{s^{\prime}}})^{-1}\circ(\exp_{p}^{g_{s^{\prime}}})^{-1}$
$=$ $(\Phi_{ss^{\prime}})^{-1}\circ(\psi_{p}^{g_{s}})^{-1}$
is a well-defined diffeomorphism of $B^{n}(r_{s}/2)$ onto its image in $M$ for all $s^{\prime}\in(s-\delta,s+\delta)\cap[0,1]$.
By (4.11), we have shown that $\exp_{p}^{g_{s^{\prime}}}:I_{p}^{g_{s^{\prime}}}(B^{n}(r_{s}/2))\to M$ is a diffeomorphism onto its image. In particular, we have proved that
$\text{\rm inj}_{g_{s^{\prime}}}(p)\geq\frac{1}{2}r_{s}$
for all $p\in M$ provided $s^{\prime}\in(s-\delta,s+\delta)\cap[0,1]$ for a sufficiently small $\delta>0$. This proves the proposition (and hence the openness of $B$) by setting $\epsilon_{s}=\frac{r_{s}}{2}$. ∎
4.2. Uniform injectivity radii lower bound near a limit point: closedness
We know that the map
$\psi_{p}^{g}:=\exp_{p}^{g}\circ I_{p}^{g}:{\mathbb{R}}^{n}\to M$ (4.12)
is well-defined by Lemma 4.2.
Remark 4.9.
The upshot of considering this map is to standardize the geodesic normal balls whose centers move around in $M$ to the fixed ball centered at the origin of ${\mathbb{R}}^{n}$. By doing so, we can consider the family $\{\psi_{p}^{g}\}_{g,p}$ whose domain is a fixed ball in the Euclidean space ${\mathbb{R}}^{n}$. These maps will be especially useful later when we try to control the injectivity radius as the centers of the geodesic balls escape to infinity.
Let $s_{i}\in B$ be a sequence converging to $s_{0}\in[0,1]$. We need to show $\iota_{g_{s_{0}}}>0$ which we will prove by contradiction. Suppose to the contrary that
$\iota_{g_{s_{0}}}=0.$ (4.13)
We start with the following obvious lemma.
Lemma 4.10.
There exists a constant $A_{0}\geq 1$ such that
$\frac{1}{A_{0}}g_{s_{0}}\leq g_{s_{i}}\leq A_{0}g_{s_{0}}$
and $\|R_{g_{s_{i}}}-R_{g_{s_{0}}}\|_{C^{0}}\to 0$ as $i\to\infty$.
At this point, there are two cases to consider:
  1. (1)
    The case where $\iota_{g_{s_{i}}}>\epsilon_{1}>0$ for all $i$,
  2. (2)
    The case where there is a subsequence, still denoted by $s_{i}$, such that $\iota_{g_{s_{i}}}\to 0$.
Case (1): An examination of the proof of Proposition 4.7 shows that the choice of $\delta>0$ therein depends only on $\epsilon>0$, the quasi-isometric ratio $A(g_{s^{\prime}},g_{s})$. Therefore we can choose $\delta>0$ independently of $i$’s by applying the same argument to all $g_{s_{i}}$ for the $\epsilon_{1}>0$ as in the proof of openness. Therefore if $i$ is sufficiently large, $s_{0}\in(s_{i}-\delta,s_{i}+\delta)$ by the standing hypothesis (4.13). This proves $\iota_{g_{s_{0}}}>0$. Therefore this case is ruled out.
Case (2): To proceed further, we will need a quantitative version of inverse function theorem or rather some idea of its proof in some circumstance. For readers’ convenience, we recall the theorem stated in [AMR83, Corollary 2.5.6] below under the $C^{2}$ bounded assumption of the relevant map $F$.
Let $g$ be any metric in the given sequence $\{g_{s_{i}}\}$ of complete Riemannian metrics on $M$ which are of bounded curvature, but a priori have no uniform lower bound away from zero for the injectivity radii.
By the curvature bound, Case (2) means that there is a pair of geodesics $\gamma_{i,p_{i}}^{\pm}(t)=\exp_{p_{i}}(tv_{i}^{\pm})$ such that
  1. (1)
    $|v_{i}^{\pm}|=1$ and $v_{i}^{+}\neq v_{i}^{-}$ contained in $T_{p_{i}}M$, and
  2. (2)
    they satisfy
    $\gamma_{i,p_{i}}^{+}(\ell_{i}^{+})=\gamma_{i,p_{i}}^{-}(\ell_{i}^{-}),$ (4.14)
    with $\ell_{i}^{\pm}\to 0$ as $i\to\infty$.
(See [CE08, Lemma 5.6] for example.) Then we have $s_{i}\to s_{0}$ and $\text{\rm inj}^{g_{s_{i}}}(p_{i})\to 0$ as $i\to\infty$.
We now quote the following quantitative inverse function theorem from [AMR83, Corollary 2.5.6] in which the theorem is stated in the general setting of Banach manifolds. (See also [Chr85, Section 8] for an essentially same statement in the finite dimensional case.)
Theorem 4.11 (Inverse Function Theorem).
Let $F$ denote an arbitrary $C^{2}$ map from a convex open ball $U\subset{\mathbb{R}}^{n}$ to ${\mathbb{R}}^{n}$, $x_{0}\in U$ and $DF(x_{0})$ is an isomorphism. We denote by $D^{n}$ and by $B^{n}$ a ball in the domain and in the codomain of the map $F$ respectively. Write
$L=\|DF(x_{0})\|,\,M=\|DF(x_{0})^{-1}\|.$
Assume $\|D^{2}F(x)\|\leq K$ for $x\in D^{n}_{R}(x_{0})$ and $\overline{D}^{n}_{R}(x_{0})\subset U$. Let
$R_{1}$ $=$ $\min\left\{\frac{1}{2KM},R\right\}$
$R_{2}$ $=$ $\min\left\{\frac{1}{R}_{1},\frac{1}{2M(L+KR_{1})}\right\},\quad R_{3}=\frac{R_{2}}{2L}.$
Then
  1. (1)
    $F$ maps the ball $D_{R_{2}}^{n}(x_{0})$ diffeomorphically onto an open set containing the ball $B_{R_{3}}^{n}(F(x_{0}))$. In particular, $F$ is one-to-one thereon.
  2. (2)
    For $y_{1},\,y_{2}\in\overline{B}^{n}_{R_{3}}(F(x_{0}))$, we have
    $\|F^{-1}(y_{1})-F^{-1}(y_{2})\|\leq 2L\|y_{1}-y_{2}\|.$
We would like to apply Theorem 4.11 to the sequence of maps given by
$F_{i}(x):=\Phi_{0s_{i}}(x)=(\psi_{p_{i}}^{g_{0}})^{-1}(\psi_{p_{i}}^{g_{s_{i}}}(x))$ (4.15)
as follows. We recall from (4.3) that the maps
$\psi_{p_{i}}^{g_{s_{i}}}=\exp_{p_{i}}^{g_{s_{i}}}\circ I_{p_{i}}^{g_{s_{i}}}$
have the full ${\mathbb{R}}^{n}$ as their domains and the full $M$ as its codomains, while the map $(\psi_{p_{i}}^{g_{0}})^{-1}$ has $B_{r_{0}}^{g_{s_{i}}}(p_{i})$ as its domain and
$U_{i}:=(\psi_{p_{i}}^{g_{0}})^{-1}(B_{r_{0}}^{g_{s_{i}}}(p_{i}))\subset{\mathbb{R}}^{n}$
as its codomain. This $U_{i}$ is an open neighborhood of $0\in{\mathbb{R}}^{n}$. Combining these, we conclude that $F_{i}$ is defined on
$(I_{p_{i}}^{g_{s_{i}}})^{-1}(B_{r_{0}}^{g_{s_{i}}}(p_{i}))\subset{\mathbb{R}}^{n}$
whose image is contained in $U_{i}\subset{\mathbb{R}}^{n}$. Here both domains and codomains are neighborhoods of $0\in{\mathbb{R}}^{n}$ respectively.
Since we have $g_{s_{i}}\to g_{s_{0}}=g$ in $C^{\infty}$ topology and $d((\psi_{p_{i}}^{g_{s_{i}}})^{-1}\circ\psi_{p_{i}}^{g_{s_{i}}})(0)=\operatorname{Id}$ by Lemma 4.8, we can find a common domain, say, $B^{n}(r_{0}/2)$ such that
$(I_{p_{i}}^{g_{s_{i}}})^{-1}(B_{r_{0}}^{g_{s_{i}}}(p_{i}))\supset B^{n}(r_{0}/2)$
and that we have
$U_{i}\subset B^{n}(r_{0})$
for all sufficiently large $i$’s: Here the constant $r_{0}>0$ is the one chosen in (4.4) depending only on the reference metric $g_{\text{\rm rf}}$ after suitably shrinking it. In summary, we have achieved the following for the maps $F_{i}$:
  • We have well-defined maps $F_{i}:B^{n}(r_{0}/2)\to B^{n}(r_{0})$ for all suffciently large $i$’s,
  • We have a uniform bound $\|D^{2}F_{i}\|<K$ for some $K>0$ independent of $i$.
This enables us to apply Theorem 4.11 to the maps $F_{i}$.
Lemma 4.12.
All $F_{i}$ satisfy the hypotheses of Theorem 4.11 with uniform constants $L,\ M$ and $R$.
Proof.
The proof is an adaptation of Cheeger’s proof of [Che70, Lemma 4.3]. We postpone the detail of the proof till Appendix A. This is another place where the bound for the derivative of the curvature and so the $C^{3}$ bound for the metrics explicitly enter. (See Lemma A.1.) ∎
This implies that $F_{i}$ is injective on $B^{n}(C^{\prime}r_{0})$ for some sufficiently small $C^{\prime}>0$ independent of $i$’s. This in turn implies that the maps $\psi_{p_{i}}^{g_{s_{i}}}$ are injective and so are $\exp_{p_{i}}^{g_{s_{i}}}$ on $B_{p_{i}}^{g_{s_{i}}}(C^{\prime}r_{0})$.
On the other hand by the standing hypothesis (4.13), the equality $\gamma_{i,p_{i}}^{+}(\ell_{i}^{+})=\gamma_{i,p_{i}}^{-}(\ell_{i}^{-})$ is equivalent to
$\exp_{p_{i}}(\ell_{i}^{+}v_{i}^{+})=\exp_{p_{i}}(\ell_{i}^{-}v_{i}^{-})$
with $\max\{\ell_{i}^{+},\ell_{i}^{-}\}<C^{\prime}r_{0}$ eventually as $i\to\infty$ which contradicts to the injectivity of $\exp_{p_{i}}$ on $B_{p_{i}}(C^{\prime}r_{0})$. Therefore Case (2) cannot occur either if we choose $r_{0}$ in $0<r_{0}<\frac{1}{4}$ sufficiently small.
Therefore we have proved $\iota_{g_{s_{0}}}>0$ by contradiction. In conclusion, we have shown $B=[0,1]$.
4.3. Uniform injectivity radii lower bound over [0,1][0,1][ 0 , 1 ]
By Proposition 4.7, for every $s\in B=[0,1]$, there exist $\delta(s),\,\epsilon(s)>0$ such that
$\iota_{g_{s^{\prime}}}>\epsilon(s)$
for all $s^{\prime}\in(s-\delta(s),s+\delta(s))\cap[0,1]$. Note that $\{(s-\delta(s),s+\delta(s))\cap[0,1]\}_{s\in[0,1]}$ is an open cover of [0,1][0,1][ 0 , 1 ]. Since [0,1][0,1][ 0 , 1 ] is compact, it has a finite open subcover, say,
$\{(s_{1}-\delta(s_{1}),s_{1}+\delta(s_{1}))\cap[0,1],\dots,(s_{m}-\delta(s_{m}),s_{m}+\delta(s_{m}))\cap[0,1]\}.$
This implies
$\iota_{g_{s}}>\min\{\epsilon(s_{1}),\dots,\epsilon(s_{m})\}>0$
for all $s\in[0,1]$. Then $\min\{\epsilon(s_{1}),\dots,\epsilon(s_{m})\}$ is the desired lower bound $\epsilon^{\prime}$ and we finished the proof of Theorem 4.3.
Remark 4.13.
It is an interesting question whether one can explicitly estimate the injectivity radius of convex sum $g_{s}$ in terms of the curvature bounds and the injectivity radii of $g_{0},\,g_{1}$. There have been studies of injectivity radii under the curvature estimates and the estimate of volume growth of geodesic balls in the literature. (See [CLY81], [CGT82] to name a couple.) These articles provide the estimates of the decay rate of injectivity radius of a point as the point diverges to infinity but no uniform lower bound away from zero. This is the reason why we have used the argument by contradiction as above since there is no mention of volume growth of $g_{0}$ or $g_{1}$ in our circumstance. In addition, there does not seem to be an easy way of formulating such a volume growth in our circumstance either, if possible at all.
5. Contractibility of the set $\operatorname{Riem}_{\mathfrak{T}}(M)$ of $C^{r}$-tame metrics for $r\geq 3$
In the present section, we will prove Theorem 1.10.
We first recall the definition of strong $C^{k}$ topology of $\operatorname{Riem}(M)$. Let $g_{\text{\rm rf}}\in{\mathfrak{T}}$ be a fixed reference metric which will be used for the study of $C^{r}$ topology on $\operatorname{Riem}_{\mathfrak{T}}(M)$. We also fix the associated Levi-Civita connection of $g_{\text{\rm rf}}\in{\mathfrak{T}}$ denoted by $\nabla$. We will denote by $D$ the associated covariant derivative applied to general tensor fields on $M$. We use $g_{\text{\rm rf}}\in{\mathfrak{T}}$ and $D$ to define the $C^{k}$-norms of general tensor fields on $M$. In particular, we will focus on the metrics in the given quasi-isometry class $\mathfrak{T}$
$g\in\operatorname{Riem}_{\mathfrak{T}}(M).$ (5.1)
A neighborhood basis element of $C^{k}$ (resp. $C^{\infty}$) topology of $\operatorname{Riem}_{\mathfrak{T}}(M)$ (with respect to $g_{\text{\rm rf}}$) at a given metric $g_{0}$ is given by the set of metrics
${\mathcal{B}}_{\{\epsilon_{i}\}}(g_{0})=\{g\in\operatorname{Riem}_{\mathfrak{T}}(M)\mid\|g-g_{0}\|<\epsilon_{0},\ldots,\|Dg\|\leq\epsilon_{1},\ldots\|D^{k}g\|<\epsilon_{r}\}$
[…]
$\Gamma(\text{\rm Sym}^{2}(TM))\supset\operatorname{Riem}_{\mathfrak{T}}(M)$
of symmetric 2-covariant tensor fields and is given by
$\|S\|:=\sup_{x\in M}|S(x)|$
for $S\in\Gamma(\text{\rm Sym}^{2}(TM))$.
Now we are ready to give the proof of the following
Theorem 5.1 (Contractibility in strong $C^{k}$ topology).
Let $r\geq 3$. Let $\mathfrak{T}$ be any quasi-isometry class of $C^{r}$-tame Riemannian metrics on $M$ and $g_{\text{\rm rf}}\in\operatorname{Riem}_{\mathfrak{T}}(M)$. Then the function
$L:[0,1]\times\operatorname{Riem}(M)\to\operatorname{Riem}(M)$ (5.2)
defined by $L(s,g)=(1-s)g+sg_{\text{\rm rf}}$ restricts to
$L:[0,1]\times\operatorname{Riem}_{\mathfrak{T}}(M)\to\operatorname{Riem}_{\mathfrak{T}}(M)$
as a continuous map, and so $\operatorname{Riem}_{\mathfrak{T}}(M)$ is contractible in strong $C^{k}$ topology for any finite $0\leq k<\infty$.
Proof.
We have already shown strict convexity of $\operatorname{Riem}_{\mathfrak{T}}(M)$ which implies the first statement. Therefore it is enough to prove the continuity of the map $L=L(s,g)$ on
$[0,1]\times\operatorname{Riem}(M).$
Here we highlight the factor $\operatorname{Riem}(M)$: the aforementioned convexity enables us to work with the factor $\operatorname{Riem}(M)$ is used for the study of continuity, not $\operatorname{Riem}_{\mathfrak{T}}(M)$, in the rest of the proof.
Let $(s_{0},g_{0})\in[0,1]\times\operatorname{Riem}(M)$, and denote by $g_{s_{0}}$ the metric given by
$g_{s_{0}}:=L(s_{0},g_{0})=(1-s_{0})g_{0}+sg_{\text{\rm rf}}$
for the simplicity of notation. Recall that any neighborhood basis element of the strong $C^{r}$ topology of $\operatorname{Riem}(M)$ at $g_{s_{0}}$ is given by the collection
=${\mathcal{B}}={\mathcal{B}}_{\{\epsilon_{i}\}}(g_{s_{0}})$
associated to the set $\{\epsilon_{i}\}=\{\epsilon_{i}\}_{0\leq i\leq r}$ of constants $\epsilon_{i}>0$. Therefore we need to prove that for each given such ${\mathcal{B}}$ the preimage
$L^{-1}({\mathcal{B}})$
is open in $[0,1]\times\operatorname{Riem}(M)$.
By the definition of product topology of $[0,1]\times\operatorname{Riem}(M)$, we need to show that there exists $\delta>0$ and a neighborhood basis element ${\mathcal{B}}_{\{\delta_{i}\}}(g_{0})$ of $g_{0}$ with $\delta_{i}$ for $i=0,\cdots,k$ such that
$L((-\delta+s_{0},\delta+s_{0})\times{\mathcal{B}}_{\{\delta_{i}\}}(g_{0}))\subset{\mathcal{B}}.$
In other words, we need to find $\delta$ and $\delta_{i}$ for $i=0,\cdots,k$ such that
$\|D^{k}(L(s,g)-L(s_{0},g_{0}))\|_{B_{\alpha}(r)}\leq\epsilon_{k},\quad g_{s_{0}}=L(s_{0},g_{0})$
for all $s\in(-\delta+s_{0},\delta+s_{0})$ and $g\in{\mathcal{B}}_{\{\delta_{i}\}}(g_{0})$. We have
$D^{k}(L(s,g)-L(s_{0},g_{0}))$ $=$ $D^{k}(L(s,g)-L(s,g_{0}))+D^{k}(L(s,g_{0})-L(s_{0},g_{0}))$
$=$ $(1-s)D^{k}(g-g_{0})+(s-s_{0})D^{k}(g_{0}).$
By the assumption from (5.1) that $g_{s_{0}}\in{\mathfrak{T}}$ and $C^{3}$-tame to $g_{\text{\rm rf}}$, we can apply Theorem 4.3 and cover $M$ by an atlas $\{B_{\alpha}(r)\}$ of balls of uniform size of radii $r=r(B_{\alpha})\geq\varepsilon>0$ for some $\varepsilon>0$. We may choose them so that $\{B_{\alpha}(r/2)\}_{\alpha}$ still covers $M$.
First we choose $\delta_{i},\,i=0,\cdots,k$ so that
$\max\big{\{}\|g-g_{0}\|_{B_{\alpha}(r)},\{\|D^{i}(g-g_{0})\|_{B_{\alpha}(r)}\}_{i=1}^{k}\big{\}}\leq\frac{1}{2}\varepsilon$
for all $\alpha$ and $g\in{\mathcal{B}}_{\{\delta_{i}\}}(g_{0})$. Then we choose $\delta>0$ so that
$\delta\max_{i=0,\cdots,k}\epsilon_{i}\leq\frac{1}{2}\varepsilon.$
Then for all $\alpha$, we obtain
$\|D^{i}(L(s,g)-L(s_{0},g_{0}))\|_{B_{\alpha}(r)}$
$\leq$ $(1-s)\|D^{i}(g-g_{0})\|_{B_{\alpha}(r)}+(s-s_{0})\|D^{i}(g_{0})\|_{B_{\alpha}(r)}$
$\leq$ $\frac{1}{2}\varepsilon+\frac{1}{2}\varepsilon=\varepsilon,$
i.e., $L(s,g)\in{\mathcal{B}}$ for all $g\in{\mathcal{B}}_{\{\delta_{i}\}}(g_{0})$ and for all $s$ with $|s-s_{0}|<\delta$. This finishes the proof of the continuity of $L$. ∎
Remark 5.2.
This proof clearly shows why it cannot be applied to the usual $C^{\infty}$ case of infinite regularity, but the proof can be adapted to the direct limit strong $C^{\infty}$ topology introduced in Appendix B.
6. No cusp-developing under the $C^{3}$-continuation of tame metrics
In this section, we prove Theorem 1.12 which we restate here.
Theorem 6.1.
Consider any continuous family $g_{s}$ in strong $C^{3}$ topology of complete Riemannian metrics with
$\|R_{g_{0}}\|_{C^{0}},\,\|DR_{g_{0}}\|_{C^{0}}<C,\quad\iota_{g_{0}}>\epsilon.$
Then there is a constant $C^{\prime}=C^{\prime}(C,\{g_{s}\}),\,\epsilon^{\prime}=\epsilon^{\prime}(C,\epsilon,\{g_{s}\})>0$ such that
$\inf_{s\in[0,1]}\iota_{g_{s}}>\epsilon^{\prime}$ (6.1)
Proof.
We first note that since $g_{s}$ is $C^{3}$-continuous and [0,1][0,1][ 0 , 1 ] is compact, the uniform curvature bound immediately follows by continuity and compactness. The quasi-isometric equivalence between $g_{0}$ and $g_{1}$ is also an immediate consequence thereof. It remains to show the injectivity radii lower bound.
Let $B\subset[0,1]$ be the subset consisting of $s$’s for which $\iota_{g_{s}}>0$. The openness of $B$ immediately follows from the curvature derivative bound and the $C^{3}$-continuity of the map $s\mapsto g_{s}$. The proof of the lower bound goes exactly the same as the proof of closedness in the previous section, which relies on the completeness of the metrics and the quantitative inverse function theorem, Theorem 4.11, especially the injectivity statement under the uniform $C^{3}$ bound.
Again we would like to apply Theorem 4.11 to the sequence of metrics $g_{s_{i}}\to g_{s_{0}}=g$ by considering the associated maps
$F_{i}:B^{n}(r_{0}/2)\to B^{n}(r_{0})$
with fixed domain and codomain given by
$F_{i}(x):=(\psi_{p}^{g_{0}})^{-1}(\psi_{p}^{g_{s_{i}}}(x))$ (6.2)
on $B^{n}(r_{0}/2)$ as before. By the given hypothesis of the derivative curvature bound and the $C^{3}$-continuity of the map $s\mapsto g_{s}$, we can again apply Lemma 4.12. This implies that $F_{i}$ is injective on $B^{n}(C^{\prime}r_{0})$ for some $C^{\prime}>0$ independent of $i$’s which again implies that $\exp_{p}^{g_{s_{i}}}$ is injective on $I_{p}^{g_{s_{i}}}(B^{n}(C^{\prime}r_{0}))\subset T_{p}M$.
Once we have achieved this far, exactly the same proof by contradiction as in the proof of Theorem 4.3, especially the closedness part thereof given in the previous Subsection 4.2 applies to derive a contradiction which in turn implies that $B=[0,1]$. This finishes the proof. ∎
Our proof of the existence of a uniform injectivity lower bound is not direct in that it does not provide an explicit bound in terms of the given geometric bound. In our proof, it is essential to consider a one-parameter family $g_{s}$ for $s\in[0,1]$ of complete metrics of uniformly bounded curvature. In this sense our result is a deformation result. The hypothesis that one of the initial metrics $g_{0}$ and $g_{1}$ has bounded curvature and injectivity lower bound away from zero, which provides a covering of the manifold by a uniform size geodesic normal coordinate charts. See [Che70] for the importance of covering the manifold.
Remark 6.2.
According to the injectivity lower bound formula from [CGT82, Equation (4.23)] it will be enough to have a lower bound for the volume growth $V_{r}^{g_{s}}(p):=\mathop{\kern 0.0pt{\rm vol}}\nolimits_{g_{s}}(B_{r}^{g_{s}}(p))$ over $s\in[0,1]$. (See [CGT82, Section 4] for some detailed discussion on the relationship between the volume lower bound and the injectivity lower bound on complete open Riemannian manifold.) It would be interesting to examine whether the uniform lower bound of the volume $V_{r}^{g_{s}}(p)$ exists for some choice of $p\in M$ and $r>0$ depending on the given one-parameter family $\{g_{s}\}$ starting from $g_{0}$ in our situation.
Part II Application to large-scale symplectic topology
7. ${\mathcal{J}}_{\omega}$ is not connected in strong $C^{\infty}$ topology
We first recall the standard definition of almost complex structures tame to a symplectic form in general as formulated by Sikorav [Sik94].
Definition 7.1.
Let $(N,J,g)$ be an almost complex manifold equipped with a Riemannian metric $g$. We say the triple is $C^{r}$-tame if the following hold:
  1. (1)
    $g$ is complete.
  2. (2)
    $g$ has bounded curvature and injectivity radius bounded away from zero. More precisely, there exist constants $C_{0}>0$ and $\varepsilon_{0}>0$ such that its curvature $R_{g}$ satisfies $\|R_{g}\|_{C^{k-2}}\leq C_{0}<\infty$ and and its injectivity radius $\iota_{g}$ satisfies $\iota_{g}\geq\varepsilon_{0}>0$.
  3. (3)
    $J$ is uniformly continuous with respect to $g$.
Condition (3) is usually rephrased into the statement that there is a tame metric $g$ and constant $C>0$ such that (1.3) holds for the pair of metrics with $(g,g_{J})$ and with $A=C$.
Remark 7.2.
  1. (1)
    It appears that Sikorav’s formulation [Sik94] starting from a Riemannian metric should be the way how one should introduce the concept of almost complex structures tame to a symplectic form in strong $C^{\infty}$ topology by remembering its quasi-isometry class of the associated metric until the end, not forgetting away along the way.
  2. (2)
    As shown in Part 1 of the present paper, we need at least the $C^{3}$-tameness to be able show that the set of tame almost complex structures is contractible. See Subsection 8.3 for the reason why.
Since the regularity is not the main issue of the present part, we will always assume that the triple $(N,J,g)$ is $C^{\infty}$-tame in the rest of the paper without further mentioning.
Recall that the notion of $J$-holomorphic curves $u$ is nothing but the same as the almost complex curves $u:(\Sigma,j)\to(N,J)$ which can be defined for any almost complex manifold.
Definition 7.3.
Let $(M,\omega)$ be a symplectic manifold. An almost complex structure $J$ on $M$ is compatible to $\omega$ if the following hold:
  1. (1)
    The bilinear form $g_{J}:=\omega(\cdot,J\cdot)$ is symmetric positive definite.
  2. (2)
    The triple $(M,J,g_{J})$ is tame in the sense of Definition of 7.1.
We denote by ${\mathcal{J}}_{\omega}$ the set of $\omega$-compatible almost complex structures and call the metric $g_{J}$ an $\omega$-tame metric.
More generally, we say $J$ is tame to $\omega$ if Condition (1) is weakened by omitting the symmetry of the bilinear form $\omega(\cdot,J\cdot)$: To any tame $\omega$, we have a canonically defined Riemannian metric given by symmetrizing the bilinear form, again denoted by $g_{J}=g_{(\omega,J)}$,
$g_{J}(v_{1},v_{2}):=\frac{\omega(v_{1},Jv_{2})+\omega(v_{2},Jv_{1})}{2}$
Recall that ${\mathcal{J}}_{\omega}$ carries the structure of an infinite dimensional smooth Fréchet manifold the tangent space of which can be written as
$T_{J}{\mathcal{J}}_{\omega}$ $=$ $\{B\in\Gamma(\text{\rm End}(E))\mid BJ+JB=0,\,\omega(B(\cdot),J(\cdot))+\omega(J(\cdot),B(\cdot))=0\}$
where the second equation means nothing but that $B$ is a symmetric endomorphism of the metric $g_{J}=\omega(\cdot,J\cdot)$ on $\xi$. (See [Flo88].)
With this preparation, Gromov’s proof of connectedness of ${\mathcal{J}}_{\omega}$ for the compact case goes as follows. By definition, ${\mathcal{J}}_{\omega}$ can be expressed as the space of smooth sections of the fiber bundle
$S_{\omega}\to M$ (7.2)
whose fiber is given by $S_{\omega,x}$ which is isomorphic to
$S({\mathbb{R}}^{2n}):=\{J_{0}\in\text{\rm Aut}({\mathbb{R}}^{2n})\mid J_{0}^{2}=-\text{\rm Id},\,-J_{0}\Omega_{0}J_{0}\,\text{\rm is positive definite}\}$
where $\Omega_{0}$ is the matrix associated to the standard symplectic bilinear form on ${\mathbb{R}}^{2n}$. When identified with ${\mathbb{R}}^{2n}\cong{\mathbb{C}}^{n}$, $\Omega_{0}$ corresponds to the complex multiplication by $\sqrt{-1}$ with the identification ${\mathbb{R}}^{2n}\cong{\mathbb{C}}^{n}$. Then this set is contractible. Gromov then concludes in [Gro85, Corollary 2.3] that ${\mathcal{J}}_{\omega}$ is contractible.
For the noncompact case, the same proof still applies in the strong $C^{3}$ topology by the result from Part 1 as long as the metric $g_{J}:=\omega(\cdot,J\cdot)$ is contained in the given quasi-isometric class of Riemannian metrics. Therefore for the Fukaya category recently constructed in [CO24] to be regarded as a symplectic invariant, one must restrict to the deformations of $J$’s that are continuous in strong $C^{r}$ topology with $3\leq r<\infty$.
Now comes one natural question arises: Are two metrics tame to the given symplectic form $\omega$ quasi-isometric? The answer to this question is simply no as the following 2 dimensional example shows.
Example 7.4.
Consider the plane $({\mathbb{R}}^{2},\omega_{0})$ with the standard symplectic form $\omega_{0}$. Consider two Riemannian metrics, $g_{0}$ the standard flat metric on ${\mathbb{R}}^{2}$ and the other the metric $g_{1}$ of the cigar
$\{(x,y,z)\mid x^{2}+y^{2}=1,\,z\geq 0\}\cup\{(x,y,z)\mid x^{2}+y^{2}+z^{2}=1,\,z\leq 0\}\subset{\mathbb{R}}^{3}$
with suitable smoothing thereof along the seam $z=0$.
Obviously they are not quasi-isometric. Both have bounded curvatures and injectivity radii bounded away from zero. Denote by $\omega_{0}=\omega_{\text{\rm flat}}$ and $\omega_{1}=\omega_{\text{\rm cigar}}$ the associated Riemannian area forms. Since both have infinite volume, there is a diffeomorphism
$\psi:({\mathbb{R}}^{2},\omega_{0})\to({\mathbb{R}}^{2},\omega_{1})$
such that $\psi^{*}\omega_{1}=\omega_{0}$ by [GS79].
Then the associated Riemannian metrics become
$\omega_{0}(\cdot,J_{0}\cdot)=g_{0},\quad\omega_{0}(\cdot,J_{1}\cdot)=g_{1},$
Therefore both $J_{0}$ and $J_{1}$ are tame to $\omega_{0}$ but their associated metrics are not quasi-isometric.
The main purpose of the present part is to identify a subset of the set ${\mathcal{J}}_{\omega}$ such that for any two $J_{1},\,J_{2}$ in the subset the associated metrics $g_{J_{1}}$ and $g_{J_{2}}$ are quasi-isometric.
We first recall the definition of strong $C^{r}$ topology of ${\mathcal{J}}_{\omega}$. (See Appendix for the definition of strong $C^{r}$ topology (resp. $C^{\infty}$ topology) on the space of sections of general fiber bundle.) Once this class of subsets is identified, the proof of disconnectedness will follow from the fact that the strong $C^{\infty}$ topology of $C^{\infty}(M,S({\mathbb{R}}^{2n}))$ is not path-connected. (See [Hir76].)
We will use the following standard notion in the comparison geometry for the study of this path-connectedness.
Definition 7.5.
For a given pair of Riemannian metrics $g_{1},\,g_{2}$, we define the function $M_{g_{1},g_{2}}:M\to{\mathbb{R}}_{+}$ given by
$M_{g_{1},g_{2}}(x)=\sup_{0\neq v\in T_{x}M}\left|\log\left(\frac{|v|_{g_{2}}}{|v|_{g_{1}}}\right)\right|$
and call the quasi-isometric ratio function of $g_{1}$ and $g_{2}$.
It follows that $M_{g_{1},g_{2}}$ is a continuous function on $M$, and its definition depends only on the $C^{0}$-values of $g_{1},\,g_{2}$ on the tangent bundle $TM$.
We now prove the following.
Proposition 7.6.
Let $(M,\omega)$ be a smooth noncompact surface equipped with an area form of infinite area, and denote by ${\mathcal{J}}_{\omega}$ the set of almost complex structures tame to $\omega$ equipped with strong $C^{\infty}$ topology. Then ${\mathcal{J}}_{\omega}$ is not path-connected.
Proof.
Denote by $\text{\rm Riem}_{\omega}(M)$ the set of $\omega$-tame Riemannian metrics
$\text{\rm Riem}_{\omega}(M)=\{g_{J}\mid J\in{\mathcal{J}}_{\omega}\}$
similarly as defined in (1.2). If ${\mathcal{J}}_{\omega}$ is path-connected, then $\text{\rm Riem}_{\omega}(M)$ also should be path-connected. We will find a pair of disjoint connected components of $\text{\rm Riem}_{\omega}(M)$ which will then show that it is not connected.
Let $g_{1},\,g_{2}$ be tame metrics. Consider the function $M_{g_{1},g_{2}}$. Now suppose that $g_{1},\,g_{2}$ are tame metrics which are not quasi-isometric. Then we can find a sequence of points $x_{1},x_{2},...$ of $M$ so that
$\lim_{n\to\infty}M_{g_{1},g_{2}}(x_{n})=\infty.$
Since two metrics on a compact set are always quasi-isometric, we may also assume that there exists a compact exhaustion $C_{1}\subset C_{2}\subset\cdots$ of $M$ so that
  • $C_{m}\subset\text{\rm Int}(C_{m+1})$ for all $m\in{\mathbb{N}}$ and
  • $\{x_{n}\}_{n\in{\mathbb{N}}}\cap C_{m}$ is finite for all $m\in{\mathbb{N}}$.
We now define an equivalence relation $\sim$ on $\text{\rm Riem}_{\omega}(M)$ by setting $g\sim h$ if and only if
$\lim_{n\to\infty}\frac{M_{g,h}(x_{n})}{M_{g_{1},g_{2}}(x_{n})}=0.$ (7.3)
We can easily check that $\sim$ is an equivalence relation $\sim$ and so its equivalence classes $\{E_{\alpha}\}_{\alpha\in I}$ for some indexing set $I$ partitions $\text{\rm Riem}_{\omega}(M)$ into a disjoint union
$\text{\rm Riem}_{\omega}(M)=\bigsqcup_{\alpha\in I}E_{\alpha}.$
We will show that each equivalence of $\sim$ is open and there are at least two nonempty equivalence classes, which will finish the proof.
Lemma 7.7.
Each equivalence class is an open subset of $\text{\rm Riem}_{\omega}(M)$.
Proof.
Pick $E_{\alpha}$ and $g\in E_{\alpha}$. We will find a strong basic open neighborhood ${\mathcal{U}}$ of $g$ such that $h\sim g$ i.e., (7.3) holds for all $h\in{\mathcal{U}}$. (see [Hir76] for the definition of strong basic open neighborhoods).
For this purpose, let $\{(\varphi_{i},U_{i})\}_{i\in\Lambda}$ be a locally finite atlas of $M$ for some countable indexing set $\Lambda$. Note that
$\{\text{\rm Int}(C_{2}),\text{\rm Int}(C_{3})\cap{C_{1}}^{c},\text{\rm Int}(C_{4})\cap{C_{2}}^{c},\dots\}$
is a locally finite open cover of $M$ and taking intersection with any locally finite atlas gives us a locally finite atlas of $M$. Moreover, the resulting atlas has additional property that the intersection between $\{x_{n}\}_{n\in{\mathbb{N}}}$ and any chart is finite. Therefore, we may assume that $K_{\ell}:=\{x_{n}\}_{n\in{\mathbb{N}}}\cap U_{\ell}$ is compact for all $\ell\in\Lambda$, and hence satisfies (7.3) for all $h\in{\mathcal{U}}$.
Furthermore by the $C^{3}$-tameness, especially by the curvature bound and the injectivity lower bound established in Part 1, we can choose the coordinate chart
$\{(U_{\ell},\varphi_{\ell})\}_{\ell\in\Lambda}$
for some indexing countable set $I$ has a uniform size with uniform bound of the second derivatives of $\varphi_{\ell}$ and its inverse.
Then a strong basic neighborhood of $g$ in $\text{\rm Riem}_{\omega}(M)$ has the form
$\{h\in\text{\rm Riem}_{\omega}(M)\mid\|\nabla^{k}h-\nabla^{k}g\|<\delta_{k},\,k=0,1,\cdots\}$
for any sequence $\delta_{k}>0$ in terms of the covariant derivatives $\nabla^{k}$ of the Levi-Civita connection of $g$, or equivalently the set with the defining inequality replaced by the norm
$\sup_{\ell\in\Lambda}\sup_{i\in{\mathbb{Z}}_{\geq 0}}\|D^{k}(h\circ\varphi_{\ell}^{-1})\|_{C^{0}(U_{i})}<\delta_{k}$
for all $k$. In particular, recalling that the definition of $M_{g,h}:M\to{\mathbb{R}}$ depends only on the $C^{0}$-values of $g,\,h$ on the tangent bundle $TM$, there exists $C=C(g;\{\delta_{k}\}_{0\leq k\leq 2})>0$ depending only on $g$ and $\{\delta_{k}\}_{0\leq k\leq 2}$ such that
$\frac{1}{C}\leq M_{g,h}(x_{n})\leq C$
[…]
We obviously have
$\lim_{n\to\infty}\frac{M_{g_{1},g_{2}}(x_{n})}{M_{g_{1},g_{2}}(x_{n})}=1\neq 0.$
which implies $g_{1}\not\sim g_{2}$. Therefore there are at least two distinct equivalence classes of $g_{1}$ and $g_{2}$. Therefore $\text{\rm Riem}_{\omega}(M)$ is not connected which finishes the proof of the proposition. ∎
Note it is well-known that the set ${\mathcal{J}}_{\omega}$ is path connected in the weak $C^{\infty}$ topology.
Remark 7.8.
There is a natural family of noncompact symplectic manifold, the symplectization of a contact manifold, more specifically the product
$(Q\times{\mathbb{R}},d(e^{s}\lambda)).$
In this case, the way how the relevant almost complex structures considered in the analysis of pseudo-holomorphic curves on symplectization is to start from CR almost complex structures $J$ on a contact manifold and consider the form of almost complex structures $\widetilde{J}=J\oplus J_{0}$ in terms of the decomposition
$T(Q\times{\mathbb{R}})=\xi\oplus{\mathcal{V}}$
where $\xi$ is the contact distribution of $Q$ and
${\mathcal{V}}:={\mathbb{R}}\left\{R_{\lambda},\frac{\partial}{\partial s}\right\}$
and $J_{0}$ is the unique almost complex structure on ${\mathcal{V}}\cong{\mathbb{R}}^{2}$ satisfying $J_{0}(\frac{\partial}{\partial s})=R_{\lambda}$. Therefore in this case there is a choice of almost complex structures that is essentially determined by the CR-almost complex structures $J$ on the given contact manifold $Q$. If $Q$ is compact, the relevant topology is the $C^{\infty}$ topology of the set of $J$’s, which also uniquely determines the relevant quasi-isometry class of the symplectization.
If $Q$ is noncompact as in the one-jet bundle one should tackle the similar topological issue for the contact manifold $Q$ in the choice of CR almost complex structures as done in [Oh21] where the notion of tame contact manifolds is introduced.
8. Quasi-isometry class $\mathfrak{T}$ and tame almost complex structures
In this section, we introduce the set of quasi-isometrically tame almost complex structures denoted by ${\mathcal{J}}_{\omega;\mathfrak{T}}$ and study its (local) contractibility.
We start with the following definition.
Definition 8.1 ($(\omega,\mathfrak{T})$-tame almost complex structures).
We call $J$ an $(\omega,\mathfrak{T})$-tame almost complex structure if the following hold:
  1. (1)
    It is $\omega$-tame.
  2. (2)
    The Riemannian metric $g_{J}=\omega(\cdot,J\cdot)$ is in the quasi-isometry class $\mathfrak{T}$.
A priori the subset ${\mathcal{J}}_{\omega;\mathfrak{T}}\subset{\mathcal{J}}_{\omega}$ is not known to be connected, because it is not known whether $g_{J}$ is tame or not for $J\in{\mathcal{J}}_{\omega}$. However we will prove that the subset ${\mathcal{J}}_{\omega;\mathfrak{T}}\subset{\mathcal{J}}_{\omega}$ is contractible in the strong $C^{r}$ topology provided $3\leq r<\infty$ for a given choice of quasi-isometry class $\mathfrak{T}$ of a metric $g$ tame to $\omega$.
Theorem 8.2.
Consider a smooth manifold $M$ and fix a quasi-isometry class $\mathfrak{T}$ of Riemannian metrics of $M$. Then the set ${\mathcal{J}}_{\omega;\mathfrak{T}}$ of quasi-isometrically tame almost complex structures is contractible with respect to the $C^{r}$ topology with $3\leq r<\infty$ of the space $\Gamma(TM)$ of sections of tangent bundles induced from that of $(M,\mathfrak{T})$.
The rest of the section will be occupied by the proof of this theorem.
In the pointwise point of view at each $x\in M$ and $t\in[0,1]$, we want to find an endomorphism $J_{x;t}:T_{x}M\to T_{x}M$ satisfying $J_{x;t}^{2}=-\operatorname{Id}$ that is compatible with $\omega_{x}$ over the family of the inner products $g_{x;t}$. Solving this linear algebra problem relies on the well-known polar decompostion in the linear algebra.
We start with recalling the polar decomposition, following the exposition of [CdS01].
8.1. Review of polar decomposition
Let $(V,\Omega)$ be a symplectic vector space. Let $G$ be any Euclidean inner product. Nondegeneracy of $\Omega$ and $G$ determine two isomorphisms
$\alpha:v\mapsto\Omega(v,\cdot)$
$\beta:v\mapsto G(v,\cdot)$
from $V$ to its dual space $V^{*}$. Then $\alpha$ is a skew-symmetric and $\beta$ is a symmetric linear map. Then we obtain a unique endomorphism $A:=\beta^{-1}\circ\alpha$ on $V$ defined by the relation
$\Omega(u,v)=G(Au,v).$ (8.1)
We denote by $A^{*}$ the adjoint linear map of $A\in\operatorname{End}(V)$ with respect to the inner product $G$. Then
$G(A^{*}u,v)=G(u,Av)=G(Av,u)=\Omega(v,u)=-\Omega(u,v)=-G(Au,v)$
and $A^{*}=-A$, so $A$ is skew-symmetric with respect to $G$. Then
  • $AA^{*}$ is symmetric (with respect to $G$): $(AA^{*})^{*}=AA^{*}$.
  • $AA^{*}$ is positive definite: $G(AA^{*}v,v)=G(A^{*}v,A^{*}v)>0$ for all $v\neq 0$.
Therefore, $AA^{*}$ is diagonalizable with positive eigenvalues $\lambda_{i}$, so that
$AA^{*}=BDB^{-1}$
for a diagonal matrix
$D=\text{\rm diag}\{\lambda_{1},\dots,\lambda_{2n}\}$ (8.2)
with $\lambda_{i}>0$. Therefore we can define the square root of $AA^{*}$ by rescaling the eigenspaces and get
$\sqrt{AA^{*}}=B\sqrt{D}B^{-1},\quad\sqrt{D}=\text{\rm diag}\{\sqrt{\lambda_{1}},\dots,\sqrt{\lambda_{2n}}\}.$
Then $\sqrt{AA^{*}}$ is again symmetric and positive definite. If we put
$J:=(\sqrt{AA^{*}})^{-1}A$ (8.3)
then $J^{2}=-\text{\rm Id}$. (The factorization $A=(\sqrt{AA^{*}})J$ is called the polar decomposition of $A$.)
The following continuity is an immediate consequence of the uniqueness of the construction.
Lemma 8.3.
Denote by $J_{G}$ the above $J$ associated to $G$ (and $\omega$). Then the map
$\text{\rm Sym}_{+}^{2}(V)\to\operatorname{End}(V);\quad G\mapsto J_{G}$
is continuous where $\text{\rm Sym}_{+}^{2}(V)$ is the set of positive definite symmetric quadratic forms on $V$.
Proof.
By definition of the matrix $A$ (8.1), we may identify $\text{\rm Sym}_{+}^{2}(V)$ with an open subset of symmetric matrices equipped with the subspace topology of the latter which is homeomorphic to ${\mathbb{R}}^{n(n+1)/2}$. Then we have the explicit expression of the map given by (8.3) from which continuity follows. ∎
8.2. Uniform pinching estimates of eigenvalues of $A_{t}A_{t}^{*}$
We fix a reference $(\omega,\mathfrak{T})$-tame almost complex structure $J_{0}\in{\mathcal{J}}_{\omega;\mathfrak{T}}$. Define $g_{0}:=\omega(\cdot,J_{0}\cdot)\in\mathfrak{T}$. By Theorem 1.9, the map
$L:[0,1]\times\text{\rm Riem}_{\mathfrak{T}}(M)\to\text{\rm Riem}_{\mathfrak{T}}(M)$
given by the linear interpolation $\gamma(t,g)=(1-t)g_{0}+tg$ is well-defined for $g\in\text{\rm Riem}_{\mathfrak{T}}(M)$ in that $g_{t}$ is contained in the same quasi-isometry class $\mathfrak{T}$.
Let $J_{1}$ be any element in ${\mathcal{J}}_{\omega;\mathfrak{T}}$ and write $g_{1}:=\omega(\cdot,J_{1}\cdot)\in\mathfrak{T}$. By definition, $g_{1}$ is $C$-quasi-isometric to $g_{0}$ for some $C\geq 1$. (Here $C=A$ in Definition 1.7.) Then $t\mapsto L(t,g_{0})$ gives us a path from $g_{0}$ to $g_{1}$. We denote
$g_{t}:=L(t,g_{0})$
which is consistent with the notation for the initial condition $L(0,g_{0})=g_{0}$. We apply polar decomposition to each $g_{t,x}$ for $(t,x)\in[0,1]\times M$ and obtain $A_{t,x}\in\operatorname{End}(T_{x}M)$. We write the associated bundle map by $A_{t}\in\operatorname{End}(TM)$.
By the unique algebraic process performed in the above polar decomposition, the assignment $t\mapsto A_{t}$ defines a continuous path in $C^{\infty}$ topology of $\operatorname{End}(TM)$. We then define
$J_{t,x}:=(\sqrt{A_{t,x}A_{t,x}^{*}})^{-1}A_{t,x}$ (8.4)
by applying (8.3) to each $t$ and $x$. We write by $J_{t}$ the associated almost complex structure of $M$.
Since $A_{t}$ commutes with $\sqrt{A_{t}A_{t}^{*}}$, $J_{t}$ commutes with $\sqrt{A_{t}A_{t}^{*}}$. Furthermore
  • $J_{t}$ is skew-symmetric with respect to $g_{t}$;
    $J_{t}^{*}=A_{t}^{*}(\sqrt{A_{t}A_{t}^{*}})^{-1}=-A_{t}(\sqrt{A_{t}A_{t}^{*}})^{-1}=-J_{t}.$
  • $J_{t}$ is orthogonal: $J_{t}^{*}J_{t}=A_{t}^{*}(\sqrt{A_{t}A_{t}^{*}})^{-1}(\sqrt{A_{t}A_{t}^{*}})^{-1}A_{t}=\text{\rm Id}$.
We check that $J_{t}$ is compatible with $\omega$: we compute
  • $\omega(J_{t}u,J_{t}v)=g_{t}(A_{t}J_{t}u,J_{t}v)=g_{t}(J_{t}A_{t}u,J_{t}v)=g_{t}(A_{t}u,v)=\omega(u,v)$
  • $\omega(v,J_{t}v)=g_{t}(A_{t}v,J_{t}v)=g_{t}(-J_{t}A_{t}v,v)=g_{t}(\sqrt{A_{t}A_{t}^{*}}v,v)>0$ for all $v\neq 0$.
Hence we obtain a $J_{t}$-tame metric $g_{J_{t}}:=\omega(\cdot,J_{t}\cdot)$ for each $t\in[0,1]$.
Now we compare the $J_{t}$-tame metric $g_{J_{t}}=\omega(\cdot,J_{t}\cdot)$ and $g_{t}$. Note that $g_{J_{t}}$ is constructed via the two-step process
$g_{t}\mapsto J_{t}\mapsto g_{J_{t}}$ (8.5)
where $g_{J_{t}}$ is not necessarily the same as the starting metric $g_{t}$. Using the definition, we compute
$g_{J_{t}}(u,v)=\omega(u,J_{t}v)=g_{t}(A_{t}u,J_{t}v)=g_{t}(J_{t}^{*}A_{t}u,v)=g_{t}(\sqrt{A_{t}A_{t}^{*}}u,v).$
Therefore we obtain the inequality
$\left(\min_{1\leq i\leq 2n}{\sqrt{\lambda_{i,t}}}\right)g_{t}(u,u)\leq g_{J_{t}}(u,u)\leq\left(\max_{1\leq i\leq 2n}{\sqrt{\lambda_{i,t}}}\right)g_{t}(u,u).$ (8.6)
This shows that once we find a lower bound (away from zero) and an upper bound of $\sqrt{\lambda_{i,t}}$, we can conclude that $g_{J_{t}}$ and $g_{t}$ are quasi-isometric.
By now, we have reduced the contractibility proof to the study of the uniform bounds for the eigenvalues $\lambda_{i,t}$ of the symmetric linear map $A_{t}A_{t}^{*}$ with respect to the metric $g_{t}$. We then examine the bounds of the eigenvalues of $\lambda_{i,t}$ henceforth.
Recall that $\omega$ is independent of $t$ and $\beta_{t}=(1-t)\beta_{0}+t\beta_{1}$ by the definition of $\beta$ since $g_{t}=(1-t)g_{0}+tg_{1}$. Now we have the linearity of the map $t\mapsto A_{t}^{-1}$:
$A_{t}^{-1}$ $=\alpha^{-1}\circ\beta_{t}=\alpha^{-1}\circ((1-t)\beta_{0}+t\beta_{1})$
$=(1-t)\alpha^{-1}\circ\beta_{0}+t\alpha^{-1}\circ\beta_{1}=(1-t)A_{0}^{-1}+tA_{1}^{-1}.$
Also note that $g_{J_{0}}=g_{0},\,g_{J_{1}}=g_{1}$ which implies $A_{0}=J_{0}$ and $A_{1}=J_{1}$. Therefore,
$A_{t}^{-1}=-(1-t)J_{0}-tJ_{1}.$
Since $A_{t}A_{t}^{*}$ is diagonalizable with positive eigenvalues, $(A_{t}A_{t}^{*})^{-1}=(A_{t}^{*})^{-1}A_{t}^{-1}$ is also diagonalizable with positive eigenvalues and
$(A_{t}^{*})^{-1}A_{t}^{-1}$ $=$ $-A_{t}^{-1}A_{t}^{-1}=-(-(1-t)J_{0}-tJ_{1})^{2}$ (8.7)
$=$ $-t^{2}J_{1}^{2}-t(1-t)J_{0}J_{1}-t(1-t)J_{1}J_{0}-(1-t)^{2}J_{0}^{2}$
$=$ $(2t^{2}-2t+1)\text{\rm{Id}}-t(1-t)J_{0}J_{1}-t(1-t)J_{1}J_{0}$
$=$ $+t(1-t)(-J_{0}J_{1}-J_{1}J_{0}-2\text{\rm{Id}}).$ Id ) .
Lemma 8.4.
We have $(-J_{1}J_{0})^{-1}=-J_{0}J_{1}$.
Proof.
We compute
$g_{0}(u,v)=\omega(u,J_{0}v)=\omega(u,-J_{1}J_{1}J_{0}v)=g_{1}(u,(-J_{1}J_{0})v),$ (8.8)
$g_{1}(u,v)=\omega(u,J_{1}v)=\omega(u,-J_{0}J_{0}J_{1}v)=g_{0}(u,(-J_{0}J_{1})v).$
The lemma immediately follows from this. ∎
[…]
Lemma 8.5.
Let $M$ be an invertible real matrix. Then every eigenvalue of $M+M^{-1}$ is of the form $\lambda_{i}+\frac{1}{\lambda_{i}}$, where $\lambda_{i}$ is an eigenvalue of $M$.
Proof.
This follows from the fact that $M$ and $M^{-1}$ commute each other and that for such a pair one can obtain simultaneous Jordan canonical forms. For completeness’ sake, we provide the details in an Appendix C. ∎
We derive
$-J_{1}J_{0}-J_{0}J_{1}=2\text{\rm{Id}}+\frac{1}{t(1-t)}((A_{t}^{*})^{-1}A_{t}^{-1}-\text{\rm{Id}})$ (8.9)
[…]
$-J_{1}J_{0}-J_{0}J_{1}=-J_{1}J_{0}+(-J_{1}J_{0})^{-1}$
from Lemma 8.4.
Now let $u_{i}$ be an eigenvector of $-J_{1}J_{0}$ with eigenvalue $\lambda_{i}$. Then $u_{i}$ is also an eigenvector of $(A_{t}^{*})^{-1}A_{t}^{-1}$ with eigenvalue
$\frac{1}{\lambda_{i,t}}:=1+t(1-t)\left(\lambda_{i}+\frac{1}{\lambda_{i}}-2\right)$ (8.10)
where $\lambda_{i,t}$ is the associated eigenvalue of $A_{t}A_{t}^{*}$.
We note that $\lambda_{i}$ is not zero and may not necessarily be real since $-J_{1}J_{0}$ is not necessarily symmetric. We also know that
  • the eigenvalue $\lambda_{i}$ gives rise to an eigenvalue of $-J_{1}J_{0}-J_{0}J_{1}$ given by $\lambda_{i}+\frac{1}{\lambda_{i}}$, and they are real by the symmetry thereof.
  • every eigenvalue of $-J_{1}J_{0}-J_{0}J_{1}$ is of this form by Lemma 8.5.
We note that the reality in the former statement is possible only when either $\lambda_{i}$ is real or it satisfies $|\lambda_{i}|=1$.
We consider the two cases separately.
Case 1: $\lambda_{i}$ is real.
If $u_{i}$ is an eigenvector of $-J_{1}J_{0}$ with real eigenvalue $\lambda_{i}$,
$g_{0}(u_{i},u_{i})=g_{1}(u_{i},-J_{1}J_{0}u_{i})=g_{1}(u_{i},\lambda_{i}u_{i})=\lambda_{i}g_{1}(u_{i},u_{i}).$
Since $g_{0}$ and $g_{1}$ are assumed to be $C$-quasi-isometric in the very beginning of the current subsection, it follows that there exists a constant $C\geq 1$ ($C=A$ in Definition 1.7) such that
$\frac{1}{C}g_{1}(u_{i},u_{i})\leq g_{0}(u_{i},u_{i})\leq Cg_{1}(u_{i},u_{i}).$
This implies $\frac{1}{C}\leq\lambda_{i}\leq C$ and we obtain
$2\leq\lambda_{i}+\frac{1}{\lambda_{i}}\leq C+\frac{1}{C}.$
We dereve the uniform bound
$1\leq\frac{1}{\lambda_{i,t}}\leq\frac{1}{2}+\frac{1}{4}\left(C+\frac{1}{C}\right)$ (8.11)
from (8.10) for all $t\in(0,1)$ holds in this case.
Case 2: $\lambda_{i}$ is not real and $|\lambda_{i}|=1$.
In this case, we have $\frac{1}{\lambda_{i}}=\overline{\lambda_{i}}$ and $\lambda_{i}+\frac{1}{\lambda_{i}}$ is a real number since the latter is an eigenvalue of the symmetric endomorphism $-J_{1}J_{0}-J_{0}J_{1}$. This in particular implies $-2<\lambda_{i}+\frac{1}{\lambda_{i}}<2$. We will now prove the following inequality.
Lemma 8.6.
Assume Case 2. Then for all $i$, we have
$\lambda_{i}+\frac{1}{\lambda_{i}}\geq 0$ (8.12)
Proof.
We prove this by contradiction.
Suppose to the contrary that there is an eigenvector $u_{i}$ of $-J_{1}J_{0}$ with eigenvalue $\lambda_{i}$ such that (8.12) fails to hold so that $\lambda_{i}+\frac{1}{\lambda_{i}}<0$. Then
$g_{0}(u_{i},(-J_{1}J_{0}-J_{0}J_{1})u_{i})=g_{0}\left(u_{i},\left(\lambda_{i}+\frac{1}{\lambda_{i}}\right)u_{i}\right)=\left(\lambda_{i}+\frac{1}{\lambda_{i}}\right)g_{0}(u_{i},u_{i})<0$
by the standing hypothesis and the positivity $g_{0}(u_{i},u_{i})>0$. On the other hand, we also have
$g_{0}(u_{i},(-J_{1}J_{0}-J_{0}J_{1})u_{i})$ $=$ $g_{0}(u_{i},-J_{1}J_{0}u_{i})+g_{0}(u_{i},-J_{0}J_{1}u_{i})$
$=$ $g_{1}(u_{i},(-J_{1}J_{0})^{2}u_{i})+g_{1}(u_{i},u_{i})$
$=$ $(\lambda_{i})^{2}g_{1}(u_{i},u_{i})+g_{1}(u_{i},u_{i})$
$=$ $((\lambda_{i})^{2}+1)g_{1}(u_{i},u_{i}).$
Therefore we obtain
$0>\left(\lambda_{i}+\frac{1}{\lambda_{i}}\right)g_{0}(u_{i},u_{i})=(\lambda_{i}^{2}+1)g_{1}(u_{i},u_{i}).$
We derive that $\lambda_{i}^{2}+1$ is also real and satisfies
$\lambda_{i}^{2}+1<0$ (8.13)
since $g_{1}(u_{i},u_{i})>0$. In particular $\lambda_{i}^{2}$ is a real number with $|\lambda_{i}|=1$, but $\lambda_{i}$ itself is not real by the standing assumption of the current case. This implies $\lambda_{i}=\pm\sqrt{-1}$, and hence $\lambda_{i}^{2}+1=0$, which contradicts to (8.13). This finishes the proof of (8.12). ∎
Corollary 8.7.
$\frac{1}{2}\leq\frac{1}{\lambda_{i,t}}<1.$ (8.14)
for all $t\in(0,1)$ in this case.
Proof.
Recall the formula (8.10)
$\frac{1}{\lambda_{i,t}}:=1+t(1-t)\left(\lambda_{i}+\frac{1}{\lambda_{i}}-2\right).$
The upper bound immediately follows which corresponds to the case $\lambda_{i}+\frac{1}{\lambda_{i}}\to 2$. On the other hand, when $\lambda_{i}+\frac{1}{\lambda_{i}}=0$, we have
$\frac{1}{\lambda_{i,t}}=1-2t(1-t)=2t^{2}-2t+1\geq\frac{1}{2}$
where the minimum is achieved at $t=\frac{1}{2}$. This finishes the proof. ∎
We summarize the above discussion into the following from the above consideration of the two cases above
Proposition 8.8.
Let $g_{0}$, $g_{1}$ and $J_{0}$, $J_{1}$ be as above, and consider the convex sum $g_{t}=(1-t)g_{0}+tg_{1}$. Then
$\frac{1}{2}\leq\frac{1}{\lambda_{i,t}}\leq\frac{1}{2}+\frac{1}{4}\left(C+\frac{1}{C}\right)$
for all $i$ and $0\leq t\leq 1$. Furthermore the constant $C$ depends only on $g_{0},\,g_{1}$.
Proof.
[…]
Remark 8.9.
  1. (1)
[…]
    $\frac{1}{t(1-t)}((A_{t}^{*})^{-1}A_{t}^{-1}-\text{\rm{Id}})$
    $=$ $-J_{1}J_{0}-J_{0}J_{1}-\frac{(1-t)}{t}J_{0}^{2}-\frac{t}{1-t}J_{1}^{2}-\frac{\text{\rm{Id}}}{t(1-t)}$
    $=$ $-J_{1}J_{0}-J_{0}J_{1}-2\text{\rm{Id}}.$
[…]
    $\frac{(1-t)}{t}+\frac{t}{1-t}-\frac{1}{t(1-t)}=-2.$
  2. (2)
    Indeed the above proof seems to display some curious linear algebraic interaction between the symplectic form and its compatible almost complex structures, which we think deserves further underpinning.
8.3. Tameness of the metric $g_{J_{t}}$
To conclude that $J_{t}$ is $(\omega,\mathfrak{T})$-tame and $J_{t}\in{\mathcal{J}}_{\omega;\mathfrak{T}}$ for $t\in(0,1)$, we have to show that $g_{J_{t}}$ is tame, i.e., complete and of bounded $\|R_{g_{J_{t}}}\|_{C^{1}}$ and has a uniform injectivity lower bound for $t\in[0,1]$. We have shown that $g_{t}$ and $g_{J_{t}}$ are quasi-isometric (and so bilipschitz) in the last subsection, and $g_{J_{t}}$ is also tame by Theorem 6.1.
9. Quasi-isometric equivalence and contractibility of ${\mathcal{J}}_{\omega;\mathfrak{T}}$
We now go back to the proof of contractibility of ${\mathcal{J}}_{\omega;\mathfrak{T}}$ and are ready to wrap-up the proof of Theorem 8.2.
[…]
We consider the map
${\mathcal{G}}:{\mathcal{J}}_{\omega;\mathfrak{T}}\to\text{\rm Riem}_{\mathfrak{T}}(M)$
given by ${\mathcal{G}}(J):=g_{J}=\omega(\cdot,J\cdot)$ which is well-defined by the definition of ${\mathcal{J}}_{\omega;\mathfrak{T}}$. We also denote by
${\mathcal{P}}:\text{\rm Riem}_{\mathfrak{T}}(M)\to{\mathcal{J}}_{\omega;\mathfrak{T}};\quad g\mapsto J_{g}$
Both maps are continuous in strong $C^{r}$ topology. The compoisition ${\mathcal{G}}\circ{\mathcal{P}}$ is nothing but the map obtained by the aforementioned two-step process (8.5) which is also continuous in strong $C^{r}$ topology.
Consider the subset
$\text{\rm Riem}_{\mathfrak{T};\omega}(M):=\{g\in\text{\rm Riem}_{\mathfrak{T}}(M)\mid g=g_{J},\,J\in{\mathcal{J}}_{\omega;\mathfrak{T}}\}.$
The following property of composition plays a fundamental role in our proof of contractibility of ${\mathcal{J}}_{\omega;\mathfrak{T}}$.
Lemma 9.1.
The image of composition ${\mathcal{G}}\circ{\mathcal{P}}$ is a retraction of
$\text{\rm Riem}_{\mathfrak{T};\omega}(M)\subset\text{\rm Riem}_{\mathfrak{T}}(M),$
i.e.,
${\mathcal{G}}\circ{\mathcal{P}}|_{\text{\rm Riem}_{\mathfrak{T};\omega}(M)}=\text{\rm Id}|_{\text{\rm Riem}_{\mathfrak{T};\omega}(M)}.$ (9.1)
Proof.
This is an immediate consequence of the uniqueness of polar decomposition: See (8.1). More specifically, we have
${\mathcal{G}}({\mathcal{P}}(g_{J}))={\mathcal{G}}(J_{g_{J}})=\omega(\cdot,J_{g_{J}}\cdot)=g_{J}$
for all $J\in{\mathcal{J}}_{\omega}$, and hence the identity (9.1). ∎
Remark 9.2.
As we mentioned before, ${\mathcal{G}}({\mathcal{P}}(g))$ is not necessarily the same as $g$ if $g\in\text{\rm Riem}_{\mathfrak{T}}(M)\setminus\text{\rm Riem}_{\mathfrak{T};\omega}(M)$.
Going back to the proof of Theorem 8.2, we denote by
${\mathcal{H}}_{g_{0}}:[0,1]\times\text{\rm Riem}_{\mathfrak{T}}(M)\to\text{\rm Riem}_{\mathfrak{T}}(M)$
the above contraction $(t,g)\mapsto(1-t)g_{0}+tg$ which is $C_{g}$-quasi-isometric by Theorem 1.9 for a constant $C_{g}\geq 1$ (depending on $g$), and hence the homotopy is well-defined and continuous in $C^{\infty}$ topology.
Then to construct the required contraction, we have only to take the composition
${\mathcal{P}}\circ{\mathcal{H}}_{g_{0}}\circ(\text{\rm Id}_{[0,1]}\times{\mathcal{G}}):[0,1]\times{\mathcal{J}}_{\omega;\mathfrak{T}}\to{\mathcal{J}}_{\omega;\mathfrak{T}}$
which defines the required contraction homotopy of ${\mathcal{J}}_{\omega;\mathfrak{T}}$. This finally finishes the proof of Theorem 8.2. ∎
The following theorem is an immediate consequence of the standard practice in symplectic topology via the machinery of pseudo-holomorphic curves.
Theorem 9.3.
Denote by $(M,\omega;\mathfrak{T})$ a (noncompact) smooth symplectic manifold $M$ with a given quasi-isometry class $\mathfrak{T}$ of Riemannian metrics on $M$. Let $\aleph(\omega,\mathfrak{T};J)$ be an invariant constructed via $J$-holomorphic curves with $J\in{\mathcal{J}}_{\omega;\mathfrak{T}}$. Then it does not depend on the choice of such $J$’s. More precisely, we have
$\aleph(\omega,\mathfrak{T};J)=\aleph(\omega,\mathfrak{T};\phi^{*}J).$
for any symplectic diffeomorphism $\phi:(M,\omega)\to(M,\omega)$ that preserves $\mathfrak{T}$, i.e., for those that satisfies $\phi^{*}J\in{\mathcal{J}}_{\omega;\mathfrak{T}}$ whenever $J\in{\mathcal{J}}_{\omega;\mathfrak{T}}$.
This leads us to consider the following subgroup of $\text{\rm Symp}(M,\omega)$ similarly as in [CO24].
Definition 9.4.
We denote by
$\text{\rm Diff}_{\text{\rm Lip}}(M,\mathfrak{T})$
the automorphism group of $\mathfrak{T}$. We then define the automorphism group of $(\omega;\mathfrak{T})$ to be the intersection
$\operatorname{Symp}(M,\omega)\cap\text{\rm Diff}_{\text{\rm Lip}}(M,\mathfrak{T})=:\operatorname{Symp}_{\text{\rm Lip}}(M,\omega;\mathfrak{T}).$
Remark 9.5.
In the present paper, we have studied the Lipschitz topology in the point of view of large-scale symplectic topology. We can also contemplate the Lipschitz topology in the point of view of micro-scale or PL topology and low regularity symplectic or contact topology. (See [KMX21] and references therein for some studies of bilipschitzian contact invariants.) We will come back to the study of the micro-scale symplectic topology elsewhere.
10. The case of 2 dimensional Riemann surfaces
In [CO24], the authors constructed a Fukaya category $Fuk(M,{\mathcal{T}})$ associated to each hyperbolic structure ${\mathcal{T}}$ on a Riemann surface, especially of infinite type. Denote by $\mathfrak{T}$ the quasi-isometry class of the hyperbolic structure ${\mathcal{T}}$. It is stated that the $A_{\infty}$ category is quasi-equivalent when one deforms almost complex structures tame to $\omega$ varies inside the set ${\mathcal{J}}({\mathcal{T}})$ consisting of ${\mathcal{T}}$-tame almost complex structures.
We fix the reference hyperbolic structure on $M$ in the sense of [LP].
Definition 10.1 (Hyperbolic Riemann surface).
A hyperbolic Riemann surface is a triple $(\Sigma,J_{0},g_{0})$ whose universal cover is isometric to the unit disk. We call it tame if it has bounded curvature and its injectivity radius is positive.
  1. (1)
    A hyperbolic structure, denoted by ${\mathcal{T}}={\mathcal{T}}_{\Sigma}$, of a surface $\Sigma$ is a choice of $(\Sigma,J_{0},g_{0}):=(\Sigma,J_{\mathcal{T}},g_{\mathcal{T}})$ that is tame.
  2. (2)
    ${\mathcal{T}}$ also determines a symplectic form
    $\omega_{\mathcal{T}}=g_{0}(J_{0}\cdot,\cdot)$
    which we call ${\mathcal{T}}$-symplectic form.
Definition 10.2 ($\omega_{\mathcal{T}}$-tame almost complex structures).
We call $J$ an $\omega_{\mathcal{T}}$-tame almost complex structure if the following holds:
  1. (1)
    It is $\omega_{\mathcal{T}}$-tame.
  2. (2)
    The metric $g_{J}=\omega_{\mathcal{T}}(\cdot,J\cdot)$ is quasi-isometric to $g_{\mathcal{T}}=\omega(\cdot,J_{\mathcal{T}}\cdot)$.
We denote by ${\mathcal{J}}({\mathcal{T}})$ the set of $\omega_{\mathcal{T}}$-tame almost complex structures.
An immediate corollary of Theorems 1.9 and 8.2 is the following special case of two-dimensions.
Theorem 10.3.
Let $\Sigma$ be a noncompact surface equipped with hyperbolic structure. Denote by $\mathfrak{T}$ be a quasi-isometry class of hyperbolic structures of $\Sigma$ and by $\operatorname{Riem}_{\mathfrak{T}}(\Sigma)$ the set of Riemannian metrics quasi-isometric to the given hyperbolic metric. Then the set $\operatorname{Riem}_{\mathfrak{T}}(\Sigma)$ is contractible in strong $C^{r}$ topology with $3\leq r<\infty$.
Remark 10.4.
This theorem in turn is used by the authors of [CO24] therein for the construction of a Fukaya category of a surface of infinite type $Fuk(\Sigma,{\mathcal{T}})$ as a quasi-isometric symplectic invariants which a priori depends on the choice of quasi-isometry class $\mathfrak{T}$ of the hyperbolic structure of the Riemann surface $\Sigma$.
Appendix A Proof of Lemma 4.12
Recall the definition
$F_{i}(x):=(\psi_{p}^{g_{0}})^{-1}(\psi_{p}^{g_{s_{i}}}(x))$
from (4.15) which are well-defined as maps
$F_{i}:B^{n}(r_{0}/2)\to B^{n}(r_{0})$
with fixed domain and codomain.
For the proof of Lemma 4.12, it is enough to prove the following estimate of the derivative of the Jacobian $\left(\frac{\partial y_{i}}{\partial x_{j}}\right)$.
Lemma A.1 (Compare with Lemma 4.3 [Che70]).
Given $S,\,S_{1}>0$, there exists a constant $C(S,S_{1})$ such that if $r_{0}<\frac{\pi}{2\sqrt{S}}$ and $\|R_{i}\|_{C^{0}}<S,\,\|DR_{i}\|_{C^{0}}<S_{1}$, then on $B^{n}(r)\subset{\mathbb{R}}^{n}$, we have
$\|D^{2}F_{i}\|_{C^{0}}\leq C(S,S_{1}).$ (A.1)
One essential difference between the framework of [Che70, Lemma 4.3] and that of Lemma A.1 is that while the former deals with coordinate change maps between two normal coordinates of the same metric, the latter involves normal coordinates of two different metrics centered at a sequence of the same point. The main interest of our study is the case when the sequence escapes to infinity on $M$. The linear isometry $I_{p}^{g}$ is used to make the center of the balls fixed at the origin in ${\mathbb{R}}^{n}$ so that we can compare the exponential maps of two different metrics $g_{0}$ and $g_{s_{i}}$.
Let $(x_{1},\cdots,x_{n})$ and $(y_{1},\cdots,y_{n})$ be normal coordinate systems of $g_{s_{i}}$ and of $g_{0}$ on $B_{r_{0}/2}(p_{i})$ and $B_{r_{0}}(p_{i})$ based on frames $\{e_{i}\}_{p_{i}}^{g_{s_{i}}},\,\{f_{i}\}_{p_{i}}^{g_{0}}$, respectively. Through the linear isometries $I_{g_{s_{i}}}^{p_{i}}$ and $I_{g_{0}}^{p_{i}}$, we may safely assume that the coordinate systems $(x_{1},\cdots,x_{n})$ and $(y_{1},\cdots,y_{n})$ are defined on ${\mathbb{R}}^{n}$ by identifying $y_{i}$ with $y_{i}\circ I_{g_{0}}^{p_{i}}$ and $x_{i}$ with $x_{i}\circ I_{g_{s_{i}}}^{p_{i}}$. After this identification, (A.1) is equivalent to the bound
$\left|\frac{\partial^{2}y_{i}}{\partial x_{k}\partial x_{j}}\right|\leq C(S,S_{1}).$ (A.2)
Then using Lemma 3.3, it will suffice to have an upper bound for the covariant derivatives
$\left\|\nabla_{\frac{\partial}{\partial x_{j}}}\right\|,\,\left\|\nabla_{\frac{\partial}{\partial y_{k}}}\right\|$
in terms of $S,\,S_{1}$. With these preparations, Cheeger’s proof of [Che70, Lemma 4.3] verbatim applies without change. This finishes the proof.
Appendix B Direct limit strong $C^{\infty}$ topology
In this appendix, we introduce the notion of direct limit strong $C^{\infty}$ topology, which is weaker that the usual definition of strong $C^{\infty}$ topology but stronger than the weak $C^{\infty}$ topology, as follows.
[…]
We consider the space $C^{\infty}({\mathcal{T}})$ of smooth sections of ${\mathcal{T}}$. Let us fix an atlas of $M$ of the type $\{\mathbb{B}_{\alpha}(r)\}$ of geodesic balls of radius $r>0$ with $r<\iota_{g}$. Then consider the filtration of $C^{\infty}({\mathcal{T}})$ given by
$C^{\infty}_{N}({\mathcal{T}}):=\{T\in\Gamma({\mathcal{T}})\mid\|D^{r}T\|<N\,\forall r\in{\mathbb{N}}\}$ (B.1)
equipped with the subspace topology of the strong $C^{\infty}$ topology. Then we have
$C^{\infty}({\mathcal{T}})=\bigcup_{N\in{\mathbb{N}}}C^{\infty}_{N}({\mathcal{T}})$
as an increasing union. We call the direct limit topology of the directed system
$C^{\infty}_{1}({\mathcal{T}})\hookrightarrow C^{\infty}_{2}({\mathcal{T}})\hookrightarrow\cdots\hookrightarrow C^{\infty}_{N}({\mathcal{T}})\cdots\longrightarrow C^{\infty}({\mathcal{T}})$
the direct limit strong $C^{\infty}$ topology. It is easy to proof the following contractibility by a slight modification of the proof of Theorem 5.1.
Theorem B.1.
The path $s\mapsto(1-s)g+sg_{\text{\rm rf}}$ is continuous with respect to the direct limit $C^{\infty}$ topology.
Appendix C Proof of Lemma 8.5
In this appendix, we give the details of the proof of Lemma 8.5.
Let $M$ be an invertible real matrix. Note that every eigenvector of $M$ is also an eigenvector of $M+M^{-1}$, but the converse might not hold in general.
Consider a Jordan decomposition of $M$. Then there exists an upper triangular matrix $U$, which is called Jordan normal form, and an invertible matrix $P$ such that
$M=PUP^{-1}.$
$U$ is a block diagonal matrix
$U=\begin{bmatrix}U_{1}&&\\ &\ddots&\\ &&U_{k}\\ \end{bmatrix}$
Each $U_{i}$ is a Jordan block of the form
$U_{i}=\begin{bmatrix}\lambda_{i}&1&&\\ &\lambda_{i}&\ddots&&\\ &&\ddots&1\\ &&&\lambda_{i}\\ \end{bmatrix}$
Where $\lambda_{i}$ is an eigenvalue of $M$ which is not zero since $M$ is invertible. Also note that
$M^{-1}=PU^{-1}P^{-1}.$
Since $U$ is an invertible block diagonal matrix,
$U^{-1}=\begin{bmatrix}U_{1}^{-1}&&\\ &\ddots&\\ &&U_{k}^{-1}\\ \end{bmatrix}$
Also, each $U_{i}$ is an upper triangular matrix and the inverse of an upper triangular matrix is upper triangular implies $U_{i}^{-1}$ is also upper triangular. Since every diagonal entry of $U_{i}$ is $\lambda_{i}$ which is not zero, every diagonal entry of $U_{i}^{-1}$ is $\frac{1}{\lambda_{i}}$.
Now we have
$M+M^{-1}$ $=PUP^{-1}+PU^{-1}P^{-1}$
$=P(U+U^{-1})P^{-1}.$
and $U+U^{-1}$ is a block diagonal matrix and each block is an upper triangular matrix whose diagonal entries are of the form $\lambda_{i}+\frac{1}{\lambda_{i}}$. Therefore, they are eigenvalues of $M+M^{-1}$. This implies that every eigenvalue of $M+M^{-1}$ is of the form $\lambda_{i}+\frac{1}{\lambda_{i}}$ and for every such eigenvalue, there exists an eigenvector $u_{i}$ of $M$ with eigenvalue $\lambda_{i}$ such that
$(M+M^{-1})u_{i}=\left(\lambda_{i}+\frac{1}{\lambda_{i}}\right)u_{i}.$
This finishes the proof.